### v1

# THERADIA WoZ: An Ecological Corpus for Appraisal-based Affect Research in Healthcare

Abstract
We present THERADIA WoZ, an ecological corpus designed for audiovisual research on affect in healthcare. Two groups of senior individuals, consisting of 52 healthy participants and 9 individuals with Mild Cognitive Impairment (MCI), performed Computerised Cognitive Training (CCT) exercises while receiving support from a virtual assistant, tele-operated by a human in the role of a Wizard-of-Oz (WoZ). The audiovisual expressions produced by the participants were fully transcribed, and partially annotated based on dimensions derived from recent models of the appraisal theories, including novelty, intrinsic pleasantness, goal conduciveness, and coping. Additionally, the annotations included 23 affective labels drew from the literature of achievement affects. We present the protocols used for the data collection, transcription, and annotation, along with a detailed analysis of the annotated dimensions and labels. Baseline methods and results for their automatic prediction are also presented. The corpus aims to serve as a valuable resource for researchers in affective computing, and is made available to both industry and academia.
Index Terms:
Ecological corpus, Computerised Cognitive Training, Appraisal theories, Dimensional / Categorical Affect Recognition
publicationid: pubid: $0$–$00$ 2021 IEEE
I Introduction
Over the past decade, special attention has been given to the use of Artificial Intelligence (AI) technologies to improve healthcare. One of the rationale behind this craze is that the automation of tasks that do not necessarily require human intervention saves therapists time and mental load, resulting in more efficient practices and less burden for both patients and clinicians. Recent reviews have however evidenced that, even though AI technologies’ performance can rival that of humans in certain assignment tasks, such as serum analysis, or assessment of cardiovascular MRI images, they still struggle to deal with social interaction skills [1, 2].
While the definition of social interaction skills remains a subject of debate [3], they can be summarised as the ability of autonomous agents to maintain equilibrium in a dynamic relationship with another agent, resulting in the development of autonomous relationships [4]. Applied to the context of interactions between AI technology and a patient in healthcare, it corresponds to the capabilities of systems to detect and respond appropriately to signals from patients that may pose a threat to the exchange equilibrium; e.g., patients may express that they feel misunderstood or unsatisfied. It is important to note that patients may convey feelings of being misunderstood through a verbal request, as well as the expression of an affective state. While verbally expressed requests are fairly well understood by today’s AI technologies, it is still difficult for them to accurately detect patients’ affective states from non-verbal signals [5].
Figure 1: Head movements, gaze, speech, and articulation of a remote operator were captured in real-time to drive a virtual assistant as a Wizard-of-Oz [6].
Yet using AI technologies for social interactions in healthcare would have far-reaching consequences. A prime illustration of their relevance lies in Computerised Cognitive Training (CCT), which aims to enhance or preserve cognitive functions in patients suffering from cognitive impairment through the repetition of exercises that target specific functions such as attention or memory [7]. Although the effectiveness of CCT has been evidenced by meta-analyses [7, 8, 9, 10], it seems to be conditioned by social interaction during the process [11], which requires the presence of the therapist. Thus, entrusting social interaction to AI technologies would be highly beneficial in making the process autonomous. To move in this direction, it is important to use meaningful data and reliable methods to accurately recognise patients’ affective state in the context of CCT.
In this study, we introduce the THERADIA WoZ corpus, an ecological corpus specifically tailored for the audiovisual detection of affective states in the healthcare domain. The corpus was built to help the development of a virtual assistant supporting CCT sessions at home, for older healthy people and patients [6]. To the best of our knowledge, this is the first audiovisual corpus of affect in healthcare that fully relies on models of appraisal theories, while being accessible to both industrial and academic research communities [^1] The URL will be provided for camera-ready.
The corpus data come from natural interactions in French involving healthy and Mild Cognitive Impairment (MCI) senior participants. The interactions revolve around CCT exercises facilitated by a virtual assistant, and operated remotely by a human acting as a Wizard-of-Oz (WoZ), cf. Figure 1. Expressions of the participants were then fully transcribed and partially annotated based on the dimensions of recent appraisal theories models, along with labels derived from the literature of achievement affects.
A comprehensive analysis of the annotated labels was carried out, enabling the identification of a core set comprising ten affective states in the context of AI technology-human interactions in healthcare. This core set of affective labels was then used along the appraisal dimensions to evaluate the performance of inference models based on audio, visual, and textual data.
The reminder of this paper is organised as follows. We review both the definition and annotation methods of affect according to the appraisal theory, along with the existing corpora in healthcare in section II. The protocol used for the construction of our data set is then presented with the material and the participants in section III. The segmentation and the annotation of the recordings are explained in section IV. A comprehensive analysis of the affective annotations is then reported in section V, and results obtained in their automatic prediction are finally reported in section VI, before concluding in section VII.
II State of the art
In this section, we first review the definition of affective states and their annotation methods through the lens of appraisal theories, and then provide a concise overview of the affect corpora available in healthcare.
II-A Affective states in the lens of the appraisal theory
According to appraisal theories, affective phenomena are defined as multi-component processes aiming to maintain the well-being and survival of individuals in their environment [12, 13]. Roughly speaking, affective phenomena would arise when the meaning of environmental stimuli are – cognitively – appraised as potentially threatening to the equilibrium of individuals in their environment.
This appraisal process would be divided into four criteria: the relevance of the events for the individual’s well-being and/or survival, their implications, the possible coping with these implications, and their normative significance, i.e., the degree of matching with individual’s norms and values [14].
Following this appraisal process, the activation of organism’s subsystems, i.e., physiological responses, action tendencies and motor expressions, would be synchronised to help individuals in carrying out the behaviors that contribute to the restoration of their homeostasis. In this perspective, an affective state serves as a marker reflecting how individuals appraise themselves between the plausible deregulation of their homeostasis caused by the environment, and the search for the behavior leading to a return to equilibrium.
The cognitive appraisal process is supposed to be causal in the elicitation of affective episodes [15]. This position assumes a necessary and sufficient relationship would link appraisal to the synchronisation of the organism’s subsystems [16]. In other words, a specific pattern of criteria of the appraisal process would trigger a specific synchronised activation of organism’s subsystems. Conversely, the observation of a specific synchronised activation would allow to predict the pattern of criteria of the appraisal process performed by the experimenter of the affective episode.
Based on these premises, the Tripartite Emotion Expression and Perception model postulates that the way humans perceive affect of others is conditioned by this causal relationship [17]. The idea is that if an individual A is able to determine the affective state of an individual B, this is because A inferred – in part unconsciously – the cognitive appraisal process performed by B on the basis of the apparent motor expressions, e.g., facial expressions, that B has produced. In brief, appraisal theories propose to solve the problem of affective state differentiation based on a $n$-dimensional hyperplane representing the criteria of the appraisal process.
II-B Differentiable annotations of affective states
II-B1 Dimensional representations
Therefore, one way of describing individuals’ affective states would be to consider them as points on a plane. Based on this approach, most of the literature have used the Circumplex model to represent them, a Cartesian plane defined by two axes, namely valence – from unpleasant to pleasant – and arousal – from deactivation to activation [18, 19]. However, this approach drew a number of criticisms for its ability to differentiate affective states; fear and anger are both defined by high arousal and negative valence, and the concept of valence itself should be rather seen as a multidimensional representation of several dimensions [20]. It has therefore been stated that more than two dimensions should be taken into account to appropriately distinguish affective states [21]. Thus, appraisal theory represents a sound framework for selecting relevant dimensions in the context of affective state annotation.
II-B2 Categorical representations
Labels are another common way of approximating individual’s affective states. One of the crucial points of such an approach is to select affective labels likely to be relevant in the context of interest. While basic emotion labels, i.e., fear, disgust, happiness, sadness, anger are commonly used to model affective states [22], they are not necessarily the most appropriate in the context of healthcare and AI-patients interactions, as they refer to affect from an evolutionary perspective.
II-B3 Relation between categorical and dimensional representations
Appraisal theories suggest that the use of labels represents a way for humans to characterise the conscious part of the activation states of the organism’s subsystems, i.e., of the cognitive appraisal process, the physiological responses, the action tendencies, and the motor expressions [23, 24], assuming that each affective label would represent a specific pattern of criteria of the appraisal process [17, 25, 26].
II-C Affect corpora for healthcare
Automatic affect analysis can be applied to various aspects of healthcare [27]: to support clinicians for the diagnosis of affective disorders [28], to facilitate the monitoring of patients’ emotions during the therapy [6], which is the focus of this study, or between therapy sessions [29].
Whereas numerous audiovisual corpora are available in the literature for affect recognition in human interactions [30], few of them were acquired in the context of healthcare. There are many reasons for this gap in the literature. One of them concerns the sensitive nature of the data that are protected by law and cannot be easily released. Moreover, populations related to healthcare are often fragile, and setting up experimental protocols to contrast affect corpora can be challenging because of their cognitive fatigue.
Nevertheless, there have been initiatives in the construction of affective corpora supporting clinicians during therapy. For instance, speech interactions between seniors and a virtual assistant playing the role of a health professional coach, were collected for the purpose of the EMPATHIC project [31], which aims to help the elderly maintain their independence as they age [32]. Authors collected $\sim$13k speech segments ($\sim$13h in total) from 136 participants over 64 years and from different countries (Spain, France, and Norway), that were further annotated by three annotators according to five labels: calm, happy, puzzled, tense, and sad, and three dimensions: arousal, valence, and dominance.
Another corpus of affect has recently been collected for children with developmental disorders using a mobile therapeutic game [33]. The corpus contains $\sim$2k video sequences of 456 children from all over the world and was annotated by 11 individuals according to seven labels: happy, surprised, sad, fearful, angry, disgust, and neutral.
III Data collection
The whole data collection protocol of this study has been documented in details, reviewed, and approved by the Ethics Committee for Research in Grenoble Alpes with the reference: CERGA-Avis-2021-1. As the experiment lasted around three hours in average, and participants were asked to consent for the commercial exploitation of their data under the protection of the European General Data Protection Regulation laws, a 20€ gift card was awarded. We also provided reimbursement for travel expanses to participants coming from a certain distance. The protocol used to collect data for the THERADIA WoZ corpus is detailed below.
III-A Participants
A total of 52 healthy older participants (40 females; mean age = 67.9, SD age = 5.1) were recruited based on advertisements published in regional newspapers, i.e., “Le Progrès” and “Le Dauphiné Libéré”. 9 participants diagnosed with MCI (two females; mean age = 75.8, SD age = 2.1) were recruited as patients of clinicians collaborating on the project. Inclusion was based on compliance with the following criteria: fluency in French, normal or corrected-to-normal vision and hearing, and consent prior to the start of the experiment. Regarding the level of education, 40% of healthy older participants had a master’s degree or higher, 33% a bachelor’s degree and 27% a baccalaureate. As for MCI participants, 22% had obtained a master’s degree or higher, 22% a bachelor’s degree, and 56% a National Vocational Qualification (French CAP), or lower.
Figure 2: Office of the experimenter operating the virtual assistant. The dialog was fully scripted and overlaid onto the participant’s video (A). The participant’s screen (B) was only visible during the exercise completion, and the virtual assistant was visible to the participant only outside the exercise completion. The virtual assistant’s rendering was consistently displayed (C) to monitor any tracking issue.
III-B CCT supported by a Virtual Assistant
All participants carried out a CCT session, which was based on eight exercises selected from the HappyNeuronPro CCT platform [^2] https://www.happyneuron.fr. The exercises involved different cognitive functions such as memory, language, attention, and planning; details on these exercises are provided in section I.A of the supplementary material.
The CCT was supported by a virtual assistant, developed by the Dynamixyz company [^3] Dynamixyz has been acquired by Take-Two Interactive in 2021., and controlled by one co-author of this study acting as a WoZ, cf. Figure 2. More specifically, head movements, gaze, speech, and articulation of the experimenter were captured in real-time to drive the 3D virtual assistant whose rendering was cast to the participant’s screen. The speech was based on a dialogue tree designed to cover a wide range of interactions, envisioned prior to the study. If none of the planned branches made sense within a given dialogue, free intervention was allowed. Eventually, the dialogue tree was updated if the unplanned interaction was deemed relevant. Technical problems that could not be solved by the operator of the virtual assistant were addressed by a second experimenter, who was sitting outside the room for the healthy participants, and inside the room for the MCI participants.
III-C Material
Questionnaires were used to interrogate various psychological, affective and cognitive dimensions of the participants, as well as to evaluate the virtual assistant and the CCT [34]. Experiments were carried out in two rooms of the EMC laboratory of Université Lyon 2, one for the participant, and one for the experimenter.
As the original lighting conditions did not allow a robust tracking of the operator of the virtual assistant, we equipped each room with an additional continuous light (Amzdeal Softbox) that provided a color temperature closed to daylight (5500K). Each room was equipped with an high-end computer, comprising a 24 inches monitor, a webcam (Logitech streamcam), a headset with a built-in unidirectional microphone (Sound blasterX H6), that was plugged to an external sound card (Presonus Audiobox), for both audio recording and streaming to the CCT application.
As pre-tests showed that the video recordings had random freezes that did not show up during the CCT session, we used an iPhone X to provide a stable video recording on the participant’s side, as well as an additional audio recording captured by a far-field omnidirectional microphone. Audio data were sampled at 44.1 kHz with 16 bits, whereas video data were sampled at a constant frame rate of 30 fps with a resolution of 1920x1080 and the YUV420p color format.
III-D Procedure
Concerning the MCI participants, the session began with the completion of the consent form and the above-mentioned questionnaires. Then, after a few instructions, the experimenter left the room so that the CCT session could start. It comprised the realisation of four exercises, each performed twice. The difficulty level of the second exercise was adjusted based on the performance of the first. It was increased in the case of success and decreased in the case of difficulty. Throughout the session, participants were guided and engaged with a virtual assistant under the impression it was operated using AI technology, whereas it was actually operated by a human as a WoZ. The story line of the session started with a welcome dialogue designed to acquaint the participant with the virtual assistant, and to collect some information such as the participant’s mood and motivation. This was followed by the exercise phase divided into five parts: the introductory dialogue detailing each exercise, the exercise completion, the feedback dialogue commenting on participant’s performance, the repetition of the exercise, as well as the feedback dialogue, and the closing dialogue, which ended the session. Once the CCT session was terminated, the whole set-up was revealed.
Concerning the healthy older participants, the procedure was the same, except for the following: First, a session consisted of eight exercises instead of four. Second, healthy older participants were divided into five groups. Twenty of them were assigned to a group without induction. The special feature of this group was that it carried out two CCT sessions, one week apart. This resulted in the set-up being revealed at the end of the second session. The remaining 32 healthy older participants were divided into four induction groups, designed to increase the chances of observing affectively-charged expressions. To this end, induction was used in one out of two of the eight exercises based on the combination of three parameters: Presenting the exercise as easy or hard beforehand; Actually setting the difficulty as easy or hard; And giving feedback from the WoZ as critical or congratulatory. Details on the groups and the inductions are reported in section I.B of the supplementary material.
TABLE I: Overall statistics of the data collected in the THERADIA-Woz.
Data
Participants
# Seq.
Seq. duration
Mean (std)
Overall
duration
Transcribed Senior 14,788 9.00 s (7.82 s) 36h59m16s
MCI 1,105 8.24 s (5.03 s) 2h31m44s
Total 15,893 8.95 s (7.66 s) 39h31m00s
Annotated Senior 2,513 8.20 s (5.27 s) 5h43m27s
MCI 222 10.09 s (6.49 s) 37m20s
Total 2,735 8.35 s (5.40 s) 6h20m47s
IV Data segmentation and annotation
We collected about 80 hours of data from the 52 healthy older partipants, and about 6 hours of data from the 9 MCI participants. We describe in this section the methods used for the segmentation, transcription, and annotation of those data.
IV-A Segmentation, transcription and selection of sequences
The segmentation of the recordings was based on variations in breath groups, ensuring they form a single proposition if they conveyed the same semantic information, or separate propositions if they related to different types of semantic information. Non-verbal expressions occurring within or proximate to a proposition were marked by a diacritic, and those occurring at distance were specifically segmented. Temporal markers for proposition start and stop were both positioned at a certain distance to allow a sufficient observation space for the affect annotation.
The transcription relied mostly on the ESLO project principles [35], respecting spelling and spoken structures; details regarding our convention are reported in section II.A of the supplementary material. Recordings were transcribed with the ELAN software [36] by five Master students in Language Engineering at the Université Grenoble Alpes with a two-steps process: an initial stage of segmentation and transcription, followed by a subsequent verification stage conducted by another student. The segmented sequences were selected to retain only those containing expressions of affect, and a fraction was randomly removed to reduce the overall duration. Obtained sequences were then all reviewed and some were further removed to reduce redundancy. Statistics of the sequences are given in table I.
IV-B Selection of affective labels and dimensions
The context of CCT aligns closely with the literature on achievement affects [37, 38, 39, 40, 41], which considers the following types of affect: (i) the activity affects – generated by the task achievement itself; (ii) the outcome affects – generated by the success or failure results, whether retrospective or prospective; (iii) the epistemic affects – generated by a form of meta-cognition on the perception of the achievement; and (iv) the social affects – generated by another individual.
These affects are particularly relevant in the context of AI-assisted healthcare, as they comprehensively cover those that can be experienced during CCT, i.e., those generated by the CCT itself (activity affects), those generated by past or future CCT failure or success (outcome affects), those generated by the perception of the cognitive abilities involved in CCT (epistemic affects), and those generated by AI-patients interactions (social affects). Our categorical representations of affect relied on the following 23 labels: angry, annoyed, anxious, ashamed, confident, contemptuous, curious, desperate, disappointed, embarrassed, excited, frustrated, interested, guilty, happy, hopeful, impatient, proud, relaxed, satisfied, sad, surprised, and upset.
The dimensions consisted of the four most important appraisal dimensions in the context of affect recognition [17, 42, 25, 26], namely: novelty, intrinsic pleasantness, goal conduciveness, and coping, plus arousal. In the context of appraisal theory, arousal is a marker of the depth of the appraisal process, which is relevant to qualify the intensity of the affective state.
IV-C Annotation tool and guidelines
We took benefits from previous collaborations with the company Viadialog [^4] https://viadialog.com/ engaged in the development of a web-based annotation tool originally tailored for audio data: ANNOT. This tool has been adapted to the needs of our project and is available online [^5] The URL will be provided for camera-ready. The platform allowed annotators – 6 French students (3 females) from different backgrounds (engineering, business, and biology) and places (Lyon and Toulouse) – to view and annotate the selected audiovisual sequences at home, on their own computer.
The annotation consisted of three steps: a time-continuous annotation, cf. Figure 3, followed by a summary value, for each of the five dimensions, and a summary value of intensity for one or multiple labels. Thus, each sequence was viewed five times, once for each dimension. At the end of the time-continuous annotation, a graph showing the dynamics of the dimension over time was presented to the annotators, who could then validate or perform the annotation again. After the validation, annotators were asked to provide a summary value for the whole sequence. Finally, once all dimensions were annotated, annotators were asked to indicate if one or more labels could describe the affect they had recognised in the sequence and provide an intensity value for them. Annotators could annotate as many labels as they wished, or none at all if the sequence did not contain any particular affect. Further details on the annotations steps are reported in section II.B of the supplementary material.
Figure 3: Example of the ANNOT web-base platform for time-continuous annotation of the intrinsic pleasantness dimension.
A document was provided for each annotator as guidelines. It was explained that the annotation task had to be performed in full screen mode in a silent place. The document advised against annotating too many sequences consecutively and suggested a ten-minute break after an hour of work. We also stressed that there were no right or wrong answers in the annotation, as the information to be judged was subjective in nature. Thus, the annotations were to be made by trying to put oneself in the participant’s shoes to assess his or her feeling about the expression produced in the selected sequences. These expressions could include the chosen words, the way they were communicated orally, and the expressive dynamics of the face and body. Finally, the document provided definitions about dimensions and labels. Each dimension was defined as follows:
  • Novelty: “To what extent does the participant feel that what is happening is predictable or unexpected?”; ranging from predictable to unexpected.
  • Intrinsic pleasantness: “At what point does the participant express pleasure or displeasure?”; ranging from unpleasantness to pleasantness.
  • Goal conduciveness: “How much does what is happening seems to correspond to the participant’s wishes?”; ranging from not at all to absolutely.
  • Coping: “How well does the participant seem able to cope with what is happening?”; ranging from not at all to absolutely.
  • Arousal: “How asleep or awake does the participant seem?”; ranging from asleep to awake.
V Corpus Analysis
This section provides a comprehensive statistical analysis of the annotation data. We first assessed the annotation frequency of the affective labels as well as the inter-annotator agreement on their frequency and intensity. We then evaluated the cognitive delay involved in the annotation of each dimension, as well as the inter-annotator agreement. Relationships between dimensions and affective labels were also estimated, and a label core set was finally selected, based on compliance with consistency criteria relating to each of the analyses mentioned above.
Figure 4: Frequency of the annotated labels according to the agreement of at least one, two or three annotators among six; the vertical axis is on a logarithmic scale to ease the reading.
V-A Categorical annotations
The frequency at which the affective labels were selected by at least one, two, or three annotators is reported in Figure 4. Based on the sample of annotation frequency per label for at least one annotator, Cohen’s $d$ was calculated to identify the theoretical value at which the effect size of its deviation from the sample mean can be considered at least "small" (i.e., Cohen’s $d$ > 0.2  [43]). This theoretical value was equal to 747 sequences. The annotation frequency per label for at least one annotator is reported in Table II, and shown in bold when above this theoretical value.
We evaluated the inter-annotator agreement on the presence of each affect label, i.e., when the intensity of a label was superior to zero, for each pair of annotators, and averaged over all pairs, by computing the Unweighted Average Recall (UAR) as follows:
$UAR=\frac{1}{k}\sum_{i=1}^{k}\frac{N_{c}^{i}}{N^{i}}$ (1)
where $k$ is the total number of classes, which corresponds here to the presence or absence of a label. $N^{i}$ is the total number of annotated sequences for a given affect label, and $N_{c}^{i}$ is the number of commonly identified classes, whether present or absent. Results are reported in Table II. According to a two-tailored $z$-test, the inter-annotator agreement would be defined as statistically above chance if it would exceed 53%, which is the case for all the labels excepted embarrassed; those labels are marked in bold style in Table II.
TABLE II: Statistics on the annotated affective labels: frequency, inter-annotator agreement on the presence, and intensity, and correlation to appraisal’s dimensions. Most relevant labels in the context of AI-assisted CCT are marked in bold.
Label Frequency Agreement on Correlation to
Presence Intensity Appraisals
Relaxed 1683 59.24 % .114
Interested 1525 54.99 % .194
Frustrated 1315 59.05 % .211
Satisfied 1216 60.69 % .097
Confident 1238 60.97 % .116
Happy 902 62.81 % .280
Surprised 851 58.44 % .260
Annoyed 856 57.52 % .130
Desperate 826 58.99 % .276
Anxious 783 54.53 % .283
Embarrassed 745 51.43 % -.256
Disappointed 520 56.00 % .387
Guilty 400 57.44 % .129
Contemptuous 363 53.61 % .359
Angry 311 56.76 % .487
Curious 328 55.60 % .439
Hopeful 223 55.62 % .480
Proud 249 58.79 % .015
Excited 159 56.39 % .150
Ashamed 144 54.22 % -.014
Impatient 138 53.14 % -.338
Upset 74 53.48 % 1.000
Sad 63 57.28 % .668
We further evaluated the inter-rater agreement on the intensity of the affective labels using the PCC. The inter-rater agreement was defined as sufficiently high if its Cohen’s $d$, i.e., its correlation coefficient value divided by the sample standard deviation, would be superior to .2 [43], which corresponds to a PCC greater than .065. This concerned most of the affective labels, which are reported in bold style in Table II.
TABLE III: Inter-annotator agreement on each sequence, or CCT session, computed with the Pearson’s correlation coefficient for both time-continuous and summary values of appraisal dimensions; mean (standard-deviation).
Sequence Session
Time-continuous values
Arousal .381 (.034) .425 (.060)
Coping .586 (.026) .573 (.032)
Goal conduciveness .602 (.019) .539 (.031)
Intrinsic pleasantness .588 (.088) .547 (.038)
Novelty .307 (.088) .294 (.128)
Summary values
Arousal .248 (.039) .228 (.065)
Coping .679 (.024) .676 (.038)
Goal conduciveness .695 (.015) .680 (.032)
Intrinsic pleasantness .662 (.020) .619 (.036)
Novelty .401 (.113) .403 (.150)
V-B Dimensional annotations
The cognitive delay appearing in the continuous annotation of each dimension was evaluated, and results showed that the delay associated with the annotation of goal conduciveness (3.3 seconds) was significantly longer than the delay observed on the other dimensions (less than 2 seconds), suggesting that annotators had more difficulty assessing it. Further details on this evaluation are provided in section III.A of the supplementary material.
We assessed the inter-annotator agreement on the dimensions for both time-continuous and summary annotations, using the Pearson Correlation Coefficient (PCC), which was computed either on each sequence, or each CCT session, cf. Table III. An high agreement was observed on the dimensions of coping, goal conduciveness, and intrinsic pleasantness, with a more pronounced consistency observed in summary annotations compared to time-continuous annotations, which is expected since time-continuous data imply a greater variation in the possible outcomes. A lower agreement was however found for arousal and novelty, which might be explained by several reasons as detailed below.
The novelty dimension was assumed to be more difficult to annotate, as it probably concerned very specific cases of our study, e.g., unplanned feedback from the virtual assistant on the participant’s performance, occurring in fewer sequences than the other dimensions. This assumption was corroborated by an analysis evidencing that the frequency of high-value annotations was lower for novelty compared to other dimensions; for a high-value threshold set at .85, the proportion of high-value annotations for novelty accounted for only .13% of the time-continuous annotation data, whereas other dimensions were at least twice more frequent: .30% for arousal, .24% for intrinsic pleasantness, .62% for goal conduciveness, and .36% for coping. A similar trend was observed in summary annotations, with the occurrence of high-value annotations for novelty accounting to .26%, whereas all other dimensions were twice as prevalent; arousal: .92%, intrinsic pleasantness: .52%, goal conduciveness: 1.35%, and coping: 2.16%.
Figure 5: Heatmap of the PCC between each dimension (summary value) presented in the columns – AR: arousal, CP: coping, GC: goal conduciveness, IP: intrinsic pleasantness, and NV: novelty – and each affective label on the rows, which are sorted in alphabetical order from positive (+) to negative (-). Significant correlations (p-value $\leq{0.05}$) are coloured either in red for positive values, or blue for negative values. Non-significant correlations (p-value $>0.05$) are presented with 0 shown in grey.
The arousal dimension was thought to be more difficult to annotate, as the context of the study, i.e. exercising on a computer and conversing with an AI, is likely to provide little variation for this dimension. This assumption is supported by the fact that the variance of arousal dimension was systematically among the lowest, which was observed for both time-continuous annotations (arousal: .010, novelty: .022, intrinsic pleasantness: .009, goal conduciveness: .017, coping: .020), and summary annotations (arousal: .008, novelty: .026, intrinsic pleasantness: .012, goal conduciveness: .021, coping: .027).
V-C Relation between dimensions and affective labels
We further analysed the agreement between annotators by exploring the relationship between the appraisal dimensions and affective labels. To measure this relationship, we calculated the PCC between each pair of dimension-affective label across all sequences and for each annotator, using the summary values. Then, we evaluated how much the annotators agreed on the correlations between affective labels and dimensions. To better understand this approach, we present as a simplified view in Figure 5 a heatmap of these associations calculated on the average of the evaluations from the six annotators (see section III.B of the supplementary material for a complete view of the heatmap of each annotator). Each cell of the heatmap corresponds to the value of the PCC calculated between a specific dimension (summary) and an affective label.
Broadly speaking, our results are in line with the predictions of appraisal theories regarding the link between affective labels and dimensions: positive affective labels are correlated with the ability to cope, goal conduciveness, and pleasantness, while negative affective labels are correlated with disability to cope, goal obstructiveness, and unpleasantness.
Some of our results are nonetheless surprising. The anger affective label is correlated with an absence of coping, whereas we would expect the opposite [44], as it generally refers to a situation in which an event is appraised as goal obstructive, but individuals appraise themselves capable of coping with it, e.g., being mugged in the street but wanting to defend themselves. In the context of CCT, where participants sit in front of a computer, different forms of anger may likely occurred, probably involving an inability to cope.
The literature suggests that the affective label of interest would be associated with a balanced mix between coping, i.e., the ability to understand, and novelty [41]. Although our results showed a correlation between interest and coping, an inverse correlation was found between interest and novelty, i.e., higher interest coincided with diminished novelty. This could be explained by the fact that the annotations in our study focused on only one sub-dimension of novelty, i.e., predictability, and did not take into account the other existing sub-dimensions such as complexity and unfamiliarity. As expected, the affective label of surprise is highly correlated with novelty. Even though the surprise affective label is generally considered to be positive or negative depending on the context, it seems to be rather negative in our results as it is correlated with unpleasantness, goal obstructiveness, and no coping. This could be explained by the fact that the occurrence of unexpected situations were more related to negative events such as unexpected negative feedback from the virtual assistant.
Although all the affective labels reported in the Figure 5 seem to be correlated with the appraisal dimensions, with the exception of sad, some correlations actually reflect the result of only one annotator. Affective labels were thus defined as irrelevant in the context AI-assisted CCT if no significant correlation is found with any appraisal dimension for at least two annotators, which was the case for several labels, as reported in Table II; labels consistent with the appraisal dimensions are marked with ✓.
Figure 6: Overview of the predictive experiments performed on the corpus.
V-D Label core set selection
A core set of the likely most relevant affective labels in the context of AI-assisted CCT was defined based on our analyses. Four main criteria of selection were used: 1) the frequency is significantly higher than the average on all affective labels, 2) the agreement on presence is significantly higher than chance, 3) the agreement on intensity is sufficiently high, and 4) there is no significant correlation with at least one dimension for at least two annotators. Ten affective labels meet these four criteria, five positive and five negative, cf. Table II.
VI Automatic recognition of affective states
In this section, we report on the automatic prediction of affective labels and dimensions. An overview of the experiments performed in this section is given in Figure 6. In brief, we extract different types of representations (expert or deep) from audio, textual, and visual data, and train specific models for the prediction of the affective labels and dimensions.
All the experiments are based on a single partitioning of the data set that was divided into three partitions: training, validation and test set, consisting of 56 %, 23 % and 21 % of the participants respectively, c.f. Table IV. These partitions were defined to ensure that the statistics – 1) ratio of female vs. male gender, 2) ratio of senior vs MCI participants, 3) ratio of participants based on affective induction, and 4) distribution of education levels – were preserved in the different population groups. All results are reported on the test set.
In the following, we introduce the used representations of the audio, visual, and textual data, along with the predictive models and strategies, and then report on the results.
TABLE IV: Summary statistics of the partitions.
Training Val Test Total
Participants 34 14 13 61
Videos 1110 851 774 2735
Duration 2h44m36s 1h51m42s 1h44m28s 6h20m47s
Population group
Senior 29 12 11 52
MCI 5 2 2 9
Gender (seniors)
Male 6 3 3 12
Female 23 9 8 40
Emotion Induction (seniors)
With induction 18 7 7 32
Without induction 11 5 4 20
Education (seniors)
High school 5 2 1 8
Professional certificate 3 1 2 6
Undergraduate 9 5 3 17
Graduate 12 4 5 21
VI-A Representations of audio, textual and video data
We included both expert and deep representations of audio, visual, and textual [^6] In order to make the comparison fair, textual data did not rely on the human transcriptions but on the outputs of an ASR, which is evaluated in details in section IV.A of the supplementary material. data to evaluate their respective contribution in the analysis of affect. Expert based features rely on the statistical description of perceptually salient patterns in the data, and involve relatively light-weight computational resources, whereas deep representations exploit self-supervised learning (SSL) methods on large scale datasets [45, 46, 47].
VI-A1 Expert based representations
We used the Mel-scale Filter Banks (MFBs), Term Frequency - Inverse Document Frequency (TF-IDF), and Facial Action Units (FAUs), as representations of the audio, textual, and visual data, respectively.
MFBs are popular features in acoustic processing. They are based on a simple signal processing pipeline which consists in taking the Fourier transform of the acoustic signal through a sliding window (25 ms shifted forward in time every 10 ms), and then mapping the power spectrum to a mel-scale (80 filter banks), which is a non-linear perceptual scale based on the human auditory system. MFBs were extracted using the SpeechBrain toolkit [48].
The TF-IDF is a classical feature in text processing, capturing the frequency of a word within a document (TF), relative to its rarity across a given set of documents (IDF). The method therefore tokenises the words of a sentence, which is represented by a vector of the count of the words it contains, and further normalised by the term frequency across the entire training partition of the dataset.
FAUs describe the intensity of movements of facial muscles related to an apparent facial expression, and are commonly used as facial descriptors in affective computing. FAUs were extracted using OpenFace [49], which provides estimations of the intensity of a set of 17 FAUs with a framerate of 30 Hz.
VI-A2 Self-supervised representations
Deep representations of audio, textual, and video data include the following models: Wav2Vec2 [50], BERT [51], and CLIP [52], respectively.
We used a multilingual Wav2Vec2 model that exploits both convolutional and transformer layers to distinguish positive and negative samples of speech in a contrastive approach [46]. The model [^7] https://huggingface.co/voidful/wav2vec2-xlsr-multilingual-56 we used has a large architecture, and was trained on 56 different languages [53], including French, and then fine-tuned for ASR.
Transformer-based representations such as BERT are popular in text processing, as they can encode the context of the words, which is crucial for affect related tasks [54, 55]. We used a BERT model [^8] https://huggingface.co/nlptown/bert-base-multilingual-uncased-sentiment fine-tuned for sentiment analysis on six languages (English, Dutch, German, French, Spanish, and Italian), and with the base architecture.
CLIP [^9] https://openai.com/research/clip is a model developed by OpenAI [52] and trained on a large dataset that includes both images and their corresponding textual descriptions allowing the model to learn the semantic relationship between visual and textual information. The architecture consists of image and text encoders that encode the input image and text into fixed-length embeddings using a contrastive learning mechanism.
VI-B Models and training strategies
The aforementioned multimodal representations were used to model the affective labels and dimensions. We evaluated the three following predictive models for the labels intensity and dimensions summaries:
  1. 1.
    Linear: One linear layer on top of the representations to map the features to the desired number of outputs.
  2. 2.
    Multi-Layer Perceptron (MLP): The linear layer as above, combined with a hidden linear layer whose size was the number of features divided by two.
  3. 3.
    Gated Recurrent Units (GRU): A GRU model with the hidden size defined to be the feature size divided by two.
Outputs of the Linear and MLP models were averaged over the sequence, whereas the last output of the GRU model was used to represent the output of the whole sequence. Preliminary results obtained in the prediction of the affective labels showed that the performance of the MLP model was consistently better than the performance of Linear and GRU models, both for the classification and regression tasks.
For dimension summaries, the model was trained to predict the five dimensions altogether in a multi-task approach. Our first experiments revealed that the Linear model consistently outperformed both the MLP and GRU models.
In order to predict the time-continuous annotations of the dimensions, we had to re-sample some features to match the annotation frame rate, which was 10 Hz. Audio representations, which had a sampling rate of 100 Hz for MFBs, and 50 Hz for Wav2vec2, were simply averaged into chunks of 10 ms to match the sampling frequency of the annotations; the same strategy was applied to the FAUs and CLIP visual representations. Regarding textual representations, we only exploited BERT as the TF-IDF returns a single feature vector for the whole sequence. Textual features were aligned to the size of the sequence using a spline interpolation, which provided a smooth and continuous representation of the features over time. As the Linear and MLP models did not achieve good performance for the prediction of time-continuous dimensions, we further experimented with various GRU models:
  1. 1.
    One-Layer GRU: A single-layered GRU with the hidden size equal to half of the input feature size, or equal to the input feature size.
  2. 2.
    Two-Layers GRU: Similar to the One-Layer GRU but with two hidden layers.
  3. 3.
    Three-Layers GRU: Similar to before, but with three hidden layers. A set of fixed hidden sizes was also tested: 128, 256, and 512.
Results showed that a GRU model with three hidden layers of the same size as the input was optimal for deep representations, whereas the same model with a hidden size of 256 yielded the best results for the hand-crafted representations.
For multimodal fusion, we experimented with middle and late fusion techniques to deal with the asynchronous inputs. The best performing approach for both labels and dimension summaries was a decision-based fusion using an Ordinary Least Squares (OLS) regression model [^10] https://scikit-learn.org. Fusion of the time-continuous dimensions was operated with a GRU as a weighted averaging layer to take into account the temporal dynamics of the features. The weights of the GRU were optimised on the training set.
All the models were trained with the Adam optimizer and an initial learning rate of $10^{-3}$ for affective labels and $10^{-4}$ for dimensions. The batch size and gradient accumulation were set to 10 for labels and 1 for dimensions. The maximum number of training epochs was set to 50 and an early stopping strategy was used with five epochs.Training of the models was done using the Mean Squared Error (MSE) as loss function, and for evaluation, the Concordance Correlation Coefficient (CCC) was calculated on all the sequences [56]. All experiments were carried out using Pytorch [^11] https://pytorch.org[57], with random seeds set to zero. The computer’s operating system was Debian GNU/Linux 10, and the GPU used to train the models was an NVIDIA Quadro RTX 6000 with 23 gigabytes of memory, CUDA version 11.3.
VI-C Affect prediction for the core set labels
Results obtained on the intensity of the labels are reported in Table V and show some strong dependencies between the modalities and the affective labels (see section III.B of the supplementary material for further details on prediction bias based on population demographics). For example, the label “happy” was more accurately evaluated by the video modality, the label “interested” by the textual modality, and the labels “frustrated” and “surprised” by the audio modality with deep representations. The choice of using either expert or deep features may further lead to important differences in the results according to the modality. For example, “frustrated” was detected with a much higher CCC (.327) with the TF-IDF compared to MFB (.151), but the opposite is observed when using the SSL representations. On average, deep representations of the textual, audio, and video modalities performed significantly better compared to their hand-crafted representations.
The fusion of the different modalities provided the best results on average, and for most affective labels. An analysis of the fusion weights showed that the relevance of each modality strongly depends on its representation. For example, when using expert features, the TF-IDF presents a much greater influence on the results compared to the MFBs and FAUs, with an averaged fusion coefficient of 0.8, compared to 0.1 for audio and video, whereas the contribution of those two modalities outweigh that of text when using deep representations, with averaged coefficients of .28, .35, and .35 for text, audio, and video respectively.
TABLE V: Core-set label intensity prediction results (CCC).
Label Expert Representations Deep Representations
Text Audio Video Multi. Text Audio Video Multi.
TF-IDF MFB FAU BERT W2V2 CLIP
Annoyed .075 .047 .274 .255 .255 .231 .311 .280
Anxious .115 .081 .129 .171 .087 .192 .123 .214
Confident .233 .176 .168 .422 .473 .482 .247 .487
Desperate .172 .148 .193 .259 .221 .277 .175 .252
Frustrated .327 .151 .175 .387 .222 .340 .239 .465
Happy .117 .273 .446 .490 .108 .303 .462 .508
Interested .147 .135 -.023 .207 .364 .204 .133 .300
Relaxed .314 .300 .385 .515 .363 .245 .440 .529
Satisfied .323 .225 .253 .428 .057 .287 .326 .486
Surprised .167 .165 .024 .197 .143 .238 .095 .275
Average .199 .170 .202 .333 .229 .280 .255 .380
An affective label was defined as present for a given sequence if at least one annotator had reported an intensity value for this label. We investigated two modeling options for their prediction: learning a binary classification model for each affective label, or learning a threshold from the regression models. As the binary classification task achieved significantly worse results compared to a decision based on the regression models, we therefore chose the latter for the experiments.
Results are shown in table VI. Similar to the previous experiments, there exists some dependencies between the best performing modality and the affective labels, even though the difference is less contrasted. Overall, we do not observe a significant difference between the expert and deep representations for the different modalities, excepted for the audio modality, and the fusion improved the performance almost systematically.
VI-D Dimensions summaries prediction
TABLE VI: Core-set label classification results (%UAR).
Label Expert Representations Deep Representations
Text Audio Video Multi. Text Audio Video Multi.
TF-IDF MFB FAU BERT W2V2 CLIP
Annoyed 63.2 57.5 62.4 66.0 63.1 63.7 65.1 69.4
Anxious 66.3 59.3 57.0 67.2 65.5 64.9 58.2 67.6
Confident 70.4 66.3 59.8 71.6 73.4 73.1 59.5 76.0
Desperate 71.3 68.6 62.4 72.1 69.3 74.9 65.6 75.7
Frustrated 72.2 63.9 63.3 74.1 70.4 69.9 63.4 74.5
Happy 65.4 62.5 71.1 73.5 65.8 66.8 73.9 76.7
Interested 61.7 57.5 50.1 62.4 64.9 60.5 59.1 64.1
Relaxed 68.7 64.9 68.7 75.1 69.0 67.3 64.1 74.0
Satisfied 70.4 63.9 64.0 72.2 73.9 70.0 63.8 75.3
Surprised 62.3 56.8 52.2 61.8 62.1 61.7 54.1 62.3
Average 67.2 62.1 61.1 69.6 67.7 67.3 62.7 71.6
Results are reported in table VII for two different gold standards: the average of annotations and the median (see section III.B of the supplementary material for further details on prediction bias based on population demographics). Comparing those two gold standards, we notice that the average of the annotations yielded the best performance across the five dimensions. Regarding the dimensions, arousal was best predicted by the audio modality and the two others modalities did not improve the performance during fusion. Arousal also achieved the lowest scores compared to the others dimensions, which might be explained by the reasons given in V-C. Coping was the best predicted dimension where all modalities performed similarly well and showed a complementarity during fusion. Goal conduciveness was best predicted by the textual representations, and fusion with the audio and video modalities also improved performance. Intrinsic pleasantness was best captured by the video modality; textual representations were relevant only when computed at the whole level of the sequence with the TF-IDF, and deep representations of audio performed similarly well. Novelty appeared to be really challenging for the video modality with extremely low scores, whereas textual representations captured better this dimension.
By analysing the relative importance of each modality based on the coefficients obtained from the linear regression model, we observed that a very low importance was given to audio features for arousal, in particular when using expert features. This may explain the decrease in performance for the multimodal approach, compared to using the audio modality alone. For the others dimensions, it was observed that the contributions of the modalities were influenced by the complexity of the representations. When using deep features, the audio modality emerged as the most influential, but this contribution was reduced for expert features, with textual information showing more important contributions.
TABLE VII: Dimension summary prediction (CCC).
Modalities
Text Audio Video Multimodal
Dimensions GS TF-IDF BERT MFB W2V2 FAU CLIP Expert Deep
Arousal Mean .311 .195 .390 .309 .234 .209 .356 .310
Median .264 .178 .276 .282 .252 .219 .326 .276
Coping Mean .506 .452 .294 .499 .336 .460 .599 .684
Median .488 .459 .302 .493 .329 .497 .603 .666
Goal conduciveness Mean .520 .495 .347 .428 .345 .380 .630 .662
Median .491 .468 .291 .426 .361 .389 .605 .642
Intrinsic pleasantness Mean .399 .313 .296 .357 .500 .573 .606 .609
Median .376 .202 .261 .346 .483 .463 .574 .594
Novelty Mean .429 .377 .357 .371 .015 .312 .465 .559
Median .329 .265 .277 .298 .021 .172 .376 .420
VI-E Continuous dimensions prediction
Results are reported in table VIII for three gold standards: the average and the median of the time-continous annotations, and a version based on the Evaluator Weighted Estimator (EWE), which corresponds to a weighted mean of the annotations using the average of the cross-CCC between each rater’s as weights [58] (see section III.B of the supplementary material for further details on prediction bias based on population demographics). Overall, the EWE approach systematically outperformed the others two gold standards on the five dimensions.
Arousal was the best predicted dimension, but this is due to the low variance observed within this dimension, as explained in the previous section. For coping, goal conduciveness and novelty, the Wav2vec2 audio representations performed better than the textual features, differently from what was observed for the summarised dimensions, which might be due to the fact that textual features provide a static representation that may not capture the continuous variations within the data. Regarding intrinsic pleasantness, the three modalities performed similarly and proved complementary during fusion, which is consistent with the observations made in the previous section. In addition, the multimodal approach consistently improved the performance. An analysis of the fusion weights revealed that when using the mean of annotations as the gold standard, the model evenly distributed importance across all modalities. However, when using the median or the EWE gold standards, the audio modality appeared as the most significant, whereas text and video contributed similarly.
An examination of the weights learned by the GRU during fusion showed that the model assigned equal importance to all modalities when using the mean of annotations as the gold standard. However, when considering the median and EWE gold standards, the audio modality emerged as the most influential, while the text and video modalities showed comparable contributions to the overall predictions.
TABLE VIII: Time-continuous dimension prediction (CCC).
Modalities
Text Audio Video Multimodal
Dimensions GS BERT MFB W2V2 FAU CLIP Deep
Arousal Mean .533 .575 .545 .539 .500 .545
Median .644 .705 .665 .696 .643 .684
EWE .554 .709 .699 .651 .643 .691
Coping Mean .356 .254 .281 .198 .239 .346
Median .186 .269 .406 .213 .260 .340
EWE .376 .308 .422 .183 .251 .418
Goal conduciveness Mean .320 .230 .236 .207 .199 .357
Median .168 .224 .390 .218 .249 .347
EWE .310 .248 .339 .191 .183 .420
Intrinsic pleasantness Mean .258 .202 .188 .262 .231 .317
Median .125 .195 .308 .272 .256 .324
EWE .262 .250 .289 .242 .245 .370
Novelty Mean .224 .154 .196 .094 .175 .231
Median .139 .116 .308 .111 .203 .274
EWE .262 .203 .279 .071 .202 .339
VII Conclusion
This paper introduced the THERADIA WoZ corpus, an annotated multimodal corpus on affect in the context of healthcare. The corpus relies on current affective science findings from recent forms of appraisal theories, and is made available to the affective computing research community. The data consist of 39.5h of multimodal sequences from 52 healthy older participants and 9 MCI participants performing WoZ-assisted CCT. Affect inductions were performed on 32 of the healthy older participants to maximise the chance of capturing affective episodes in the data. All sequences were fully transcribed and partially annotated based on four appraisal dimensions (i.e., novelty, intrinsic pleasantness, goal conduciveness, and coping) plus the arousal dimension, as well as 23 achievement affect labels. Dimensions were annotated both in continuous time and in summarised form, while labels were only annotated in summarised form in terms of presence and intensity. Combining the results of different analyses yields a core set of ten affective labels (five positive, five negative), evidenced as highly relevant in the context of AI-assisted CCT.
Extensive baseline experiments and analyses were then conducted to model affect for both labels and dimensions using three modalities of text, audio and video. We also explored the fusion of different modalities for both computationally inexpensive feature extraction techniques and more contextual deep representations. Results showed some dependencies between the modalities and the different representations of affect. This work provides new insights into the field of affect recognition in healthcare and helps to provide additional data for benchmark comparisons.
In conclusion, we believe that this multidisciplinary work, combining recent advances in affective computing and psychology, fits perfectly with current research challenges. Indeed, we believe that such collaborations will benefit both affective computing and psychology by pushing back the boundaries of our knowledge of affect and how it can be modelled.
VIII Acknowledgments
This research has received funding from the Banque Publique d’Investissement (BPI) under grant agreement THERADIA, the Association Nationale de la Recherche et de la Technologie (ANRT), under grant agreement No. 2019/0729, and has been partially supported by MIAI@Grenoble-Alpes, (ANR-19-P3IA-0003).

### v2

@@
 The audiovisual expressions produced by the participants were fully transcribed, and partially annotated based on dimensions derived from recent models of the appraisal theories, including novelty, intrinsic pleasantness, goal conduciveness, and coping.
-Additionally, the annotations included 23 affective labels drew from the literature of achievement affects.
+Additionally, the annotations included 23 affective labels drawn from the literature of achievement affects.
 We present the protocols used for the data collection, transcription, and annotation, along with a detailed analysis of the annotated dimensions and labels.
@@
 Over the past decade, special attention has been given to the use of Artificial Intelligence (AI) technologies to improve healthcare.
-One of the rationale behind this craze is that the automation of tasks that do not necessarily require human intervention saves therapists time and mental load, resulting in more efficient practices and less burden for both patients and clinicians.
-Recent reviews have however evidenced that, even though AI technologies’ performance can rival that of humans in certain assignment tasks, such as serum analysis, or assessment of cardiovascular MRI images, they still struggle to deal with social interaction skills [1, 2].
+One of the rationales behind this craze is that the automation of tasks that do not necessarily require human intervention saves therapists time and mental load, resulting in more efficient practices and less burden for both patients and clinicians.
+However, recent reviews have evidenced that, even though AI technologies’ performance can rival that of humans in certain specific tasks, such as serum analysis, or assessment of cardiovascular MRI images, they still struggle to deal with social interaction skills [1].
 While the definition of social interaction skills remains a subject of debate [2], they can be summarised as the ability of autonomous agents to maintain equilibrium in a dynamic relationship with another agent, resulting in the development of autonomous relationships [3].
-Applied to the context of interactions between AI technology and a patient in healthcare, it corresponds to the capabilities of systems to detect and respond appropriately to signals from patients that may pose a threat to the exchange equilibrium; e.g., patients may express that they feel misunderstood or unsatisfied.
-It is important to note that patients may convey feelings of being misunderstood through a verbal request, as well as the expression of an affective state.
+In the context of interactions between AI technology and patients in healthcare, it corresponds to the capabilities of systems to detect and respond appropriately to signals from patients that may disrupt the balance of interaction; e.g., patients may express that they feel misunderstood or unsatisfied.
 While verbally expressed requests are fairly well understood by today’s AI technologies, it is still difficult for them to accurately detect patients’ affective states from non-verbal signals [4].
 Figure 1: Head movements, gaze, speech, and articulation of a remote operator were captured in real-time to drive a virtual assistant as a Wizard-of-Oz [5].
-Yet using AI technologies for social interactions in healthcare would have far-reaching consequences.
+Yet, the use of AI technologies for social interactions in healthcare would have far-reaching consequences.
 A prime illustration of their relevance lies in Computerised Cognitive Training (CCT), which aims to enhance or preserve cognitive functions in patients suffering from cognitive impairment through the repetition of exercises that target specific functions such as attention or memory [6].
-Although the effectiveness of CCT has been evidenced by meta-analyses [7, 8, 9, 10], it seems to be conditioned by social interaction during the process [11], which requires the presence of the therapist.
+Although the effectiveness of CCT has been evidenced by meta-analyses [6, 7], it seems to be conditioned by social interaction during the process [8], which requires the presence of the therapist.
 Thus, entrusting social interaction to AI technologies would be highly beneficial in making the process autonomous.
 To move in this direction, it is important to use meaningful data and reliable methods to accurately recognise patients’ affective state in the context of CCT.
-In this study, we introduce the THERADIA WoZ corpus, an ecological corpus specifically tailored for the audiovisual detection of affective states in the healthcare domain.
+In this study, we present the THERADIA WoZ corpus, an ecological corpus specifically tailored for the audiovisual detection of affective states in the healthcare domain.
 The corpus was built to help the development of a virtual assistant supporting CCT sessions at home, for older healthy people and patients [5].
@@
 A comprehensive analysis of the annotated labels was carried out, enabling the identification of a core set comprising ten affective states in the context of AI technology-human interactions in healthcare.
-This core set of affective labels was then used along the appraisal dimensions to evaluate the performance of inference models based on audio, visual, and textual data.
-The reminder of this paper is organised as follows.
-We review both the definition and annotation methods of affect according to the appraisal theory, along with the existing corpora in healthcare in section II.
+This core set of affective labels was then used along with the appraisal dimensions to evaluate the performance of inference models based on audio, visual, and textual data.
+The remainder of this paper is organised as follows.
+We first review the appraisal theory and annotation methods of affect, along with the existing corpora in healthcare in section II.
 The protocol used for the construction of our data set is then presented with the material and the participants in section III.
@@
 II State of the art
-In this section, we first review the definition of affective states and their annotation methods through the lens of appraisal theories, and then provide a concise overview of the affect corpora available in healthcare.
+In this section, we first review the definition of affective states and their annotation methods through the lens of appraisal theories, and then provide a concise overview of corpora of affect available in healthcare.
 II-A Affective states in the lens of the appraisal theory
-According to appraisal theories, affective phenomena are defined as multi-component processes aiming to maintain the well-being and survival of individuals in their environment [12, 13].
-Roughly speaking, affective phenomena would arise when the meaning of environmental stimuli are – cognitively – appraised as potentially threatening to the equilibrium of individuals in their environment.
+According to appraisal theories, affective phenomena are defined as multi-component processes aiming to maintain the well-being and survival of individuals in their environment [9].
+Roughly speaking, affective phenomena arise when the meaning of environmental stimuli are cognitively appraised as potentially threatening to an individual’s equilibrium in their environment.
 This appraisal process would be divided into four criteria: the relevance of the events for the individual’s well-being and/or survival, their implications, the possible coping with these implications, and their normative significance, i.e., the degree of matching with individual’s norms and values [10].
-Following this appraisal process, the activation of organism’s subsystems, i.e., physiological responses, action tendencies and motor expressions, would be synchronised to help individuals in carrying out the behaviors that contribute to the restoration of their homeostasis.
-In this perspective, an affective state serves as a marker reflecting how individuals appraise themselves between the plausible deregulation of their homeostasis caused by the environment, and the search for the behavior leading to a return to equilibrium.
+Following this appraisal process, the activation of the organism’s subsystems, i.e., physiological responses, action tendencies and motor expressions, would be synchronised to help individuals carry out the behaviours that contribute to the restoration of their homeostasis.
+In this perspective, an affective state serves as a marker reflecting how individuals appraise themselves between the plausible deregulation of their homeostasis caused by the environment, and the search for the behaviour leading to a return to equilibrium.
 The cognitive appraisal process is supposed to be causal in the elicitation of affective episodes [11].
 This position assumes a necessary and sufficient relationship would link appraisal to the synchronisation of the organism’s subsystems [12].
-In other words, a specific pattern of criteria of the appraisal process would trigger a specific synchronised activation of organism’s subsystems.
+In other words, a specific pattern of criteria of the appraisal process would trigger a specific synchronised activation of the organism’s subsystems.
 Conversely, the observation of a specific synchronised activation would allow to predict the pattern of criteria of the appraisal process performed by the experimenter of the affective episode.
-Based on these premises, the Tripartite Emotion Expression and Perception model postulates that the way humans perceive affect of others is conditioned by this causal relationship [17].
+Based on these premises, the Tripartite Emotion Expression and Perception model postulates that the way humans perceive the affect of others is conditioned by this causal relationship [13].
 The idea is that if an individual $A$ is able to determine the affective state of an individual $B$, this is because $A$ inferred – in part unconsciously – the cognitive appraisal process performed by $B$ on the basis of the apparent motor expressions, e.g., facial expressions, that $B$ has produced.
@@
 Therefore, one way of describing individuals’ affective states would be to consider them as points on a plane.
-Based on this approach, most of the literature have used the Circumplex model to represent them, a Cartesian plane defined by two axes, namely valence – from unpleasant to pleasant – and arousal – from deactivation to activation [18, 19].
+Based on this approach, most of the literature have used the Circumplex model to represent them, a Cartesian plane defined by two axes, namely valence – from unpleasant to pleasant – and arousal – from deactivation to activation [14].
 However, this approach drew a number of criticisms for its ability to differentiate affective states; fear and anger are both defined by high arousal and negative valence, and the concept of valence itself should be rather seen as a multidimensional representation of several dimensions [15].
@@
 One of the crucial points of such an approach is to select affective labels likely to be relevant in the context of interest.
-While basic emotion labels, i.e., fear, disgust, happiness, sadness, anger are commonly used to model affective states [22], they are not necessarily the most appropriate in the context of healthcare and AI-patients interactions, as they refer to affect from an evolutionary perspective.
+While basic emotion labels, i.e., fear, disgust, happiness, sadness, and anger are commonly used to model affective states [17], they are not necessarily the most appropriate in the context of healthcare and AI-patients interactions, as they refer to affect from an evolutionary perspective.
 II-B3 Relation between categorical and dimensional representations
-Appraisal theories suggest that the use of labels represents a way for humans to characterise the conscious part of the activation states of the organism’s subsystems, i.e., of the cognitive appraisal process, the physiological responses, the action tendencies, and the motor expressions [23, 24], assuming that each affective label would represent a specific pattern of criteria of the appraisal process [17, 25, 26].
+Appraisal theories suggest that the use of labels represents a way for humans to characterise the conscious part of the activation states of the organism’s subsystems, i.e., of the cognitive appraisal process, the physiological responses, the action tendencies, and the motor expressions [18], assuming that each affective label would represent a specific pattern of criteria of the appraisal process [13, 19].
 II-C Affect corpora for healthcare
-Automatic affect analysis can be applied to various aspects of healthcare [27]: to support clinicians for the diagnosis of affective disorders [28], to facilitate the monitoring of patients’ emotions during the therapy [6], which is the focus of this study, or between therapy sessions [29].
+Automatic affect analysis can be applied to various aspects of healthcare [20]: to support clinicians for the diagnosis of affective disorders, to facilitate the monitoring of patients’ emotions during the therapy [5], which is the focus of this study, or between therapy sessions.
 Whereas numerous audiovisual corpora are available in the literature for affect recognition in human interactions [21], few of them were acquired in the context of healthcare.
-There are many reasons for this gap in the literature.
+There are several reasons for this gap in the literature.
 One of them concerns the sensitive nature of the data that are protected by law and cannot be easily released.
@@
 Nevertheless, there have been initiatives in the construction of affective corpora supporting clinicians during therapy.
-For instance, speech interactions between seniors and a virtual assistant playing the role of a health professional coach, were collected for the purpose of the EMPATHIC project [31], which aims to help the elderly maintain their independence as they age [32].
+For instance, speech interactions between seniors and a virtual assistant playing the role of a health professional coach, were collected for the purpose of the EMPATHIC project [22], which aims to help the elderly maintain their independence as they age.
 Authors collected $\sim$13k speech segments ($\sim$13h in total) from 136 participants over 64 years and from different countries (Spain, France, and Norway), that were further annotated by three annotators according to five labels: calm, happy, puzzled, tense, and sad, and three dimensions: arousal, valence, and dominance.
@@
 III Data collection
-The whole data collection protocol of this study has been documented in details, reviewed, and approved by the Ethics Committee for Research in Grenoble Alpes with the reference: CERGA-Avis-2021-1.
-As the experiment lasted around three hours in average, and participants were asked to consent for the commercial exploitation of their data under the protection of the European General Data Protection Regulation laws, a 20€ gift card was awarded.
-We also provided reimbursement for travel expanses to participants coming from a certain distance.
+The whole data collection protocol of this study has been documented in detail, reviewed, and approved by the Ethics Committee for Research in Grenoble Alpes with the reference: CERGA-Avis-2021-1.
+As the experiment lasted around three hours on average, and participants were asked to consent for the commercial exploitation of their data under the protection of the European General Data Protection Regulation laws, a 20€ gift card was awarded.
+Participants could ask for reimbursement of travel expenses, i.e., including public transport fares or petrol costs, for both the journey to and from the experiment site.
 The protocol used to collect data for the THERADIA WoZ corpus is detailed below.
 III-A Participants
-A total of 52 healthy older participants (40 females; mean age = 67.9, SD age = 5.1) were recruited based on advertisements published in regional newspapers, i.e., “Le Progrès” and “Le Dauphiné Libéré”.
-9 participants diagnosed with MCI (two females; mean age = 75.8, SD age = 2.1) were recruited as patients of clinicians collaborating on the project.
-Inclusion was based on compliance with the following criteria: fluency in French, normal or corrected-to-normal vision and hearing, and consent prior to the start of the experiment.
-Regarding the level of education, 40% of healthy older participants had a master’s degree or higher, 33% a bachelor’s degree and 27% a baccalaureate.
+A total of 52 healthy senior participants (40 females; mean age = 67.9, SD age = 5.1) were recruited based on advertisements published in regional newspapers such as “Le Progrès” and “Le Dauphiné Libéré”.
+Additionnaly, 9 participants diagnosed with MCI (two females; mean age = 75.8, SD age = 2.1) were recruited via clinicians collaborating on the project.
+Inclusion criteria included fluency in French, either normal or corrected-to-normal vision and hearing, and informed consent prior to recordings.
+Regarding the level of education, 40% of healthy senior participants had a master’s degree or higher, 33% a bachelor’s degree and 27% a baccalaureate.
 As for MCI participants, 22% had obtained a master’s degree or higher, 22% a bachelor’s degree, and 56% a National Vocational Qualification (French CAP), or lower.
@@
 III-B CCT supported by a Virtual Assistant
-All participants carried out a CCT session, which was based on eight exercises selected from the HappyNeuronPro CCT platform [^2] https://www.happyneuron.fr.
+All participants carried out a CCT session, which was based on eight exercises selected from the HappyNeuronPro CCT platform.
 The exercises involved different cognitive functions such as memory, language, attention, and planning; details on these exercises are provided in section I.A of the supplementary material.
-The CCT was supported by a virtual assistant, developed by the Dynamixyz company [^3] Dynamixyz has been acquired by Take-Two Interactive in 2021., and controlled by one co-author of this study acting as a WoZ, cf.
+The CCT was supported by a virtual assistant, developed by the Dynamixyz company, and controlled by one co-author of this study acting as a WoZ, cf.
 Figure 2.
@@
 Questionnaires were used to interrogate various psychological, affective and cognitive dimensions of the participants, as well as to evaluate the virtual assistant and the CCT [24].
-Experiments were carried out in two rooms of the EMC laboratory of Université Lyon 2, one for the participant, and one for the experimenter.
+Experiments were carried out in two quiet rooms of the EMC laboratory of Université Lyon 2, one for the participant, and one for the experimenter.
 As the original lighting conditions did not allow a robust tracking of the operator of the virtual assistant, we equipped each room with an additional continuous light (Amzdeal Softbox) that provided a color temperature closed to daylight (5500K).
 Each room was equipped with an high-end computer, comprising a 24 inches monitor, a webcam (Logitech streamcam), a headset with a built-in unidirectional microphone (Sound blasterX H6), that was plugged to an external sound card (Presonus Audiobox), for both audio recording and streaming to the CCT application.
-As pre-tests showed that the video recordings had random freezes that did not show up during the CCT session, we used an iPhone X to provide a stable video recording on the participant’s side, as well as an additional audio recording captured by a far-field omnidirectional microphone.
+As pre-tests revealed random freezes in the video recordings that were not apparent during the CCT session, we used an iPhone X to provide a stable video recording on the participant’s side, along with an additional audio recording captured by a far-field omnidirectional microphone.
 Audio data were sampled at 44.1 kHz with 16 bits, whereas video data were sampled at a constant frame rate of 30 fps with a resolution of 1920x1080 and the YUV420p color format.
@@
 It was increased in the case of success and decreased in the case of difficulty.
-Throughout the session, participants were guided and engaged with a virtual assistant under the impression it was operated using AI technology, whereas it was actually operated by a human as a WoZ.
+Throughout the session, participants were guided and engaged with the virtual assistant under the impression it was operated using AI technology, whereas it was actually operated by a human as a WoZ.
 The story line of the session started with a welcome dialogue designed to acquaint the participant with the virtual assistant, and to collect some information such as the participant’s mood and motivation.
@@
 Once the CCT session was terminated, the whole set-up was revealed.
-Concerning the healthy older participants, the procedure was the same, except for the following: First, a session consisted of eight exercises instead of four.
-Second, healthy older participants were divided into five groups.
+Concerning the healthy senior participants, the procedure was the same, except for the following: First, a session consisted of eight exercises instead of four.
+Second, healthy senior participants were divided into five groups.
 Twenty of them were assigned to a group without induction.
@@
 This resulted in the set-up being revealed at the end of the second session.
-The remaining 32 healthy older participants were divided into four induction groups, designed to increase the chances of observing affectively-charged expressions.
-To this end, induction was used in one out of two of the eight exercises based on the combination of three parameters: Presenting the exercise as easy or hard beforehand; Actually setting the difficulty as easy or hard; And giving feedback from the WoZ as critical or congratulatory.
+The remaining 32 healthy senior participants were divided into four induction groups, designed to increase the chances of observing affectively-charged expressions.
+To this end, induction was used in one out of two of the eight exercises based on the combination of three parameters: (i) presenting the exercise as easy or hard beforehand; (ii) actually setting the difficulty as easy or hard; and (iii) giving feedback from the WoZ as critical or congratulatory.
 Details on the groups and the inductions are reported in section I.B of the supplementary material.
-TABLE I: Overall statistics of the data collected in the THERADIA-Woz.
+TABLE I: Overall statistics of the data collected in the THERADIA-Woz
 Data
@@
 We collected about 80 hours of data from the 52 healthy older partipants, and about 6 hours of data from the 9 MCI participants.
-We describe in this section the methods used for the segmentation, transcription, and annotation of those data.
+We describe in this section the methods used for the segmentation, transcription, and annotation of this data.
 IV-A Segmentation, transcription and selection of sequences
-The segmentation of the recordings was based on variations in breath groups, ensuring they form a single proposition if they conveyed the same semantic information, or separate propositions if they related to different types of semantic information.
+The segmentation of the recordings was based on the pragmatic completeness in breath groups, i.e., groups of words that are delimited by pauses taken for breathing, ensuring they formed a single proposition if they conveyed the same semantic information, or separate propositions if they related to different types of semantic information [25].
 Non-verbal expressions occurring within or proximate to a proposition were marked by a diacritic, and those occurring at distance were specifically segmented.
-Temporal markers for proposition start and stop were both positioned at a certain distance to allow a sufficient observation space for the affect annotation.
-The transcription relied mostly on the ESLO project principles [35], respecting spelling and spoken structures; details regarding our convention are reported in section II.A of the supplementary material.
+Temporal markers for proposition start and stop were both positioned at a certain distance to allow a sufficient observation space for the annotation of affect.
+The transcription relied mostly on the ESLO project principles [26], respecting spelling and spoken structures; details regarding our conventions are reported in section II.A of the supplementary material.
 Recordings were transcribed with the ELAN software [27] by five Master students in Language Engineering at the Université Grenoble Alpes with a two-steps process: an initial stage of segmentation and transcription, followed by a subsequent verification stage conducted by another student.
@@
 IV-B Selection of affective labels and dimensions
-The context of CCT aligns closely with the literature on achievement affects [37, 38, 39, 40, 41], which considers the following types of affect: (i) the activity affects – generated by the task achievement itself; (ii) the outcome affects – generated by the success or failure results, whether retrospective or prospective; (iii) the epistemic affects – generated by a form of meta-cognition on the perception of the achievement; and (iv) the social affects – generated by another individual.
+The context of CCT aligns closely with the literature on achievement affects [28, 29, 30], which considers the following types of affect:
+  1. (i)
+    activity affects generated by the task achievement itself;
+  2. (ii)
+    outcome affects generated by the success or failure results, whether retrospective or prospective;
+  3. (iii)
+    epistemic affects generated by a form of metacognition on the perception of the achievement;
+  4. (iv)
+    social affects generated by another individual.
 These affects are particularly relevant in the context of AI-assisted healthcare, as they comprehensively cover those that can be experienced during CCT, i.e., those generated by the CCT itself (activity affects), those generated by past or future CCT failure or success (outcome affects), those generated by the perception of the cognitive abilities involved in CCT (epistemic affects), and those generated by AI-patients interactions (social affects).
-Our categorical representations of affect relied on the following 23 labels: angry, annoyed, anxious, ashamed, confident, contemptuous, curious, desperate, disappointed, embarrassed, excited, frustrated, interested, guilty, happy, hopeful, impatient, proud, relaxed, satisfied, sad, surprised, and upset.
+Our categorical representations of affect relied on the following 23 labels: angry, annoyed, anxious, ashamed, confident, contemptuous, curious, desperate, disappointed, embarrassed, excited, frustrated, interested, guilty, happy, hopeful, impatient, proud, relaxed, sad, satisfied, surprised, and upset.
 The dimensions consisted of the four most important appraisal dimensions in the context of affect recognition [13, 31, 32, 19], namely: novelty, intrinsic pleasantness, goal conduciveness, and coping, plus arousal.
@@
 IV-C Annotation tool and guidelines
-We took benefits from previous collaborations with the company Viadialog [^4] https://viadialog.com/ engaged in the development of a web-based annotation tool originally tailored for audio data: ANNOT.
+We took benefit from previous collaborations with the company Viadialog [^2] https://viadialog.com/ engaged in the development of a web-based annotation tool originally tailored for audio data: ANNOT.
 This tool has been adapted to the needs of our project and is available online [^3] The URL will be provided for camera-ready.
-The platform allowed annotators – 6 French students (3 females) from different backgrounds (engineering, business, and biology) and places (Lyon and Toulouse) – to view and annotate the selected audiovisual sequences at home, on their own computer.
+The platform allowed annotators – here 6 French students (3 females) from different backgrounds (engineering, business, and biology) and places (Lyon and Toulouse) – to view and annotate the selected audiovisual sequences at home, using their own computer.
 The annotation consisted of three steps: a time-continuous annotation, cf.
@@
 Finally, once all dimensions were annotated, annotators were asked to indicate if one or more labels could describe the affect they had recognised in the sequence and provide an intensity value for them.
-Annotators could annotate as many labels as they wished, or none at all if the sequence did not contain any particular affect.
-Further details on the annotations steps are reported in section II.B of the supplementary material.
+Annotators could annotate as many labels as desired, or none at all if the sequence lacked any particular affect.
+Further details on the annotations steps are reported in section II.C of the supplementary material.
 Figure 3: Example of the ANNOT web-base platform for time-continuous annotation of the intrinsic pleasantness dimension.
-A document was provided for each annotator as guidelines.
-It was explained that the annotation task had to be performed in full screen mode in a silent place.
-The document advised against annotating too many sequences consecutively and suggested a ten-minute break after an hour of work.
-We also stressed that there were no right or wrong answers in the annotation, as the information to be judged was subjective in nature.
-Thus, the annotations were to be made by trying to put oneself in the participant’s shoes to assess his or her feeling about the expression produced in the selected sequences.
-These expressions could include the chosen words, the way they were communicated orally, and the expressive dynamics of the face and body.
-Finally, the document provided definitions about dimensions and labels.
+Before starting, annotators participated in a training session with the project investigators, including an expert in appraisal theories of emotion.
+The purpose of this session was to familiarise the annotators with the concepts of appraisal theories, i.e., the meaning of dimensions and affective labels, and to explain in detail the guidelines for the annotation process.
+Annotators were instructed to put themselves in the participant’s shoes to assess the emotional expression conveyed in the selected sequences, considering factors such as the chosen words, their oral delivery, and the expressive dynamics of the face and body.
+It was emphasised that there were no right or wrong answers, as the task was inherently subjective.
+Additionally, they were instructed to perform the task in full-screen mode, in a quiet environment, with sessions lasting no longer than one hour, followed by ten-minute breaks.
+The annotation process was continuously monitored through an online tracking table, updated daily by the annotators, with additional communication via email or phone as needed.
+After the training session, each annotator received a document summarising all these points, including definitions of dimensions and labels.
 Each dimension was defined as follows:
@@
 This section provides a comprehensive statistical analysis of the annotation data.
-We first assessed the annotation frequency of the affective labels as well as the inter-annotator agreement on their frequency and intensity.
+We first assessed the annotation frequency of the affective labels as well as the inter-annotator agreement on their intensity.
 We then evaluated the cognitive delay involved in the annotation of each dimension, as well as the inter-annotator agreement.
 Relationships between dimensions and affective labels were also estimated, and a label core set was finally selected, based on compliance with consistency criteria relating to each of the analyses mentioned above.
-Figure 4: Frequency of the annotated labels according to the agreement of at least one, two or three annotators among six; the vertical axis is on a logarithmic scale to ease the reading.
+Figure 4: Frequency (logarithmic scale) of the annotated labels according to the agreement of at least one, two or three annotators among six.
 V-A Categorical annotations
 The frequency at which the affective labels were selected by at least one, two, or three annotators is reported in Figure 4.
-Based on the sample of annotation frequency per label for at least one annotator, Cohen’s $d$ was calculated to identify the theoretical value at which the effect size of its deviation from the sample mean can be considered at least "small" (i.e., Cohen’s $d$ > 0.2  [43]).
-This theoretical value was equal to 747 sequences.
+Based on the sample of annotation frequency per label for at least one annotator, Cohen’s $d$ was calculated to identify the theoretical value at which the effect size of its deviation from the sample mean can be considered at least "small", i.e., Cohen’s $d$ > 0.2  [33]; this theoretical value was equal to 747 sequences.
 The annotation frequency per label for at least one annotator is reported in Table II, and shown in bold when above this theoretical value.
@@
 Results are reported in Table II.
-According to a two-tailored $z$-test, the inter-annotator agreement would be defined as statistically above chance if it would exceed 53%, which is the case for all the labels excepted embarrassed; those labels are marked in bold style in Table II.
+According to a two-tailored $z$-test, the inter-annotator agreement would be defined as statistically above chance if it would exceed 53%, which is the case for all the labels excepted embarrassed.
 TABLE II: Statistics on the annotated affective labels: frequency, inter-annotator agreement on the presence, and intensity, and correlation to appraisal’s dimensions.
-Most relevant labels in the context of AI-assisted CCT are marked in bold.
-Label Frequency Agreement on Correlation to
+Most relevant labels in the context of AI-assisted CCT are marked in bold
+  Label Frequency Agreement on Cor. to
 Presence Intensity Appraisals
@@
 Impatient 138 53.14 % -.338
-Upset 74 53.48 % 1.000
+Upset 74 53.48 % 1.00
 Sad 63 57.28 % .668
-We further evaluated the inter-rater agreement on the intensity of the affective labels using the PCC.
-The inter-rater agreement was defined as sufficiently high if its Cohen’s $d$, i.e., its correlation coefficient value divided by the sample standard deviation, would be superior to .2 [43], which corresponds to a PCC greater than .065.
-This concerned most of the affective labels, which are reported in bold style in Table II.
-TABLE III: Inter-annotator agreement on each sequence, or CCT session, computed with the Pearson’s correlation coefficient for both time-continuous and summary values of appraisal dimensions; mean (standard-deviation).
+We further evaluated the inter-rater agreement on the intensity of the affective labels using the Pearson’s Correlation Coefficient (PCC).
+The inter-rater agreement was defined as sufficiently high if its Cohen’s $d$, i.e., its PCC value divided by the sample standard deviation, is superior to .2 [33], which corresponds to a PCC greater than .065.
+This concerned most of the affective labels, which are reported in bold style.
+TABLE III: Inter-annotator agreement on each sequence, or CCT session, computed with the Pearson’s correlation coefficient for both time-continuous and summary values of appraisal dimensions; mean (standard-deviation)
 Sequence Session
@@
 V-B Dimensional annotations
-The cognitive delay appearing in the continuous annotation of each dimension was evaluated, and results showed that the delay associated with the annotation of goal conduciveness (3.3 seconds) was significantly longer than the delay observed on the other dimensions (less than 2 seconds), suggesting that annotators had more difficulty assessing it.
-Further details on this evaluation are provided in section III.A of the supplementary material.
-We assessed the inter-annotator agreement on the dimensions for both time-continuous and summary annotations, using the Pearson Correlation Coefficient (PCC), which was computed either on each sequence, or each CCT session, cf.
+The cognitive delay in the continuous annotation of each dimension was evaluated, and results showed that goal conduciveness (3.3 seconds) had a significantly longer delay than the other dimensions (less than 2 seconds), suggesting that annotators had more difficulty assessing it.
+This evaluation is detailed in section III.A of the supplementary material.
+We assessed the inter-annotator agreement on the dimensions for both time-continuous and summary annotations, using the PCC, which was computed either on each sequence, or each CCT session, cf.
 Table III.
@@
 The novelty dimension was assumed to be more difficult to annotate, as it probably concerned very specific cases of our study, e.g., unplanned feedback from the virtual assistant on the participant’s performance, occurring in fewer sequences than the other dimensions.
-This assumption was corroborated by an analysis evidencing that the frequency of high-value annotations was lower for novelty compared to other dimensions; for a high-value threshold set at .85, the proportion of high-value annotations for novelty accounted for only .13% of the time-continuous annotation data, whereas other dimensions were at least twice more frequent: .30% for arousal, .24% for intrinsic pleasantness, .62% for goal conduciveness, and .36% for coping.
+This assumption was corroborated by an analysis evidencing that the frequency of high-value annotations was lower for novelty compared to other dimensions; for a high-value threshold set at .85, the proportion of high-value annotations for novelty accounted for only .13% of the time-continuous annotation data, whereas other dimensions were at least twice as frequent: .30% for arousal, .24% for intrinsic pleasantness, .62% for goal conduciveness, and .36% for coping.
 A similar trend was observed in summary annotations, with the occurrence of high-value annotations for novelty accounting to .26%, whereas all other dimensions were twice as prevalent; arousal: .92%, intrinsic pleasantness: .52%, goal conduciveness: 1.35%, and coping: 2.16%.
-Figure 5: Heatmap of the PCC between each dimension (summary value) presented in the columns – AR: arousal, CP: coping, GC: goal conduciveness, IP: intrinsic pleasantness, and NV: novelty – and each affective label on the rows, which are sorted in alphabetical order from positive (+) to negative (-).
-Significant correlations (p-value $\leq{0.05}$) are coloured either in red for positive values, or blue for negative values.
-Non-significant correlations (p-value $>0.05$) are presented with 0 shown in grey.
+Figure 5: Heatmap of the PCC between affective labels and summary values of appraisal dimensions aggregated across all annotators.
+The sign (+) or (-) after each label denotes whether it is positive or negative, respectively.
+Dimensions are as follows: AR: arousal, CP: coping, GC: goal conduciveness, IP: intrinsic pleasantness, NV: novelty.
+The correlations shown represent the average between annotators for whom the correlation was significant (p < 0.05) after Bonferroni correction and of identical direction.
+Positive or negative direction is shown in blue or red, respectively.
+A correlation present for a single annotator was considered null.
 The arousal dimension was thought to be more difficult to annotate, as the context of the study, i.e.
-exercising on a computer and conversing with an AI, is likely to provide little variation for this dimension.
-This assumption is supported by the fact that the variance of arousal dimension was systematically among the lowest, which was observed for both time-continuous annotations (arousal: .010, novelty: .022, intrinsic pleasantness: .009, goal conduciveness: .017, coping: .020), and summary annotations (arousal: .008, novelty: .026, intrinsic pleasantness: .012, goal conduciveness: .021, coping: .027).
+exercising on a computer and conversing with a virtual assistant, is likely to provide little variation for this dimension.
+This assumption is supported by the fact that the variance of arousal dimension was systematically among the lowest, which was observed for both time-continuous annotations – arousal: .010, novelty: .022, intrinsic pleasantness: .009, goal conduciveness: .017, coping: .020 –, and summary annotations – arousal: .008, novelty: .026, intrinsic pleasantness: .012, goal conduciveness: .021, coping: .027.
 V-C Relation between dimensions and affective labels
-We further analysed the agreement between annotators by exploring the relationship between the appraisal dimensions and affective labels.
-To measure this relationship, we calculated the PCC between each pair of dimension-affective label across all sequences and for each annotator, using the summary values.
-Then, we evaluated how much the annotators agreed on the correlations between affective labels and dimensions.
-To better understand this approach, we present as a simplified view in Figure 5 a heatmap of these associations calculated on the average of the evaluations from the six annotators (see section III.B of the supplementary material for a complete view of the heatmap of each annotator).
-Each cell of the heatmap corresponds to the value of the PCC calculated between a specific dimension (summary) and an affective label.
+Further analysis of annotator agreement was conducted by exploring the relationship between summary values of appraisal dimensions and affective labels of each annotator separately.
+Specifically, the PCC was calculated between each pair of appraisal dimensions and affective labels for all sequences in which the value of the label concerned differed from zero; cf.
+section III.B of the supplementary material for a comprehensive view of the heatmap of each annotator.
+A correlation was considered consistent if it was statistically significant and in the same direction for at least two annotators.
+The average of consistent correlations between annotators are reported in Figure 5.
+The labels that are consistently correlated with at least one appraisal dimension are marked with "✓" in the Table II.
 Broadly speaking, our results are in line with the predictions of appraisal theories regarding the link between affective labels and dimensions: positive affective labels are correlated with the ability to cope, goal conduciveness, and pleasantness, while negative affective labels are correlated with disability to cope, goal obstructiveness, and unpleasantness.
-Some of our results are nonetheless surprising.
+Some of our results are, however, unexpected.
 The anger affective label is correlated with an absence of coping, whereas we would expect the opposite [34], as it generally refers to a situation in which an event is appraised as goal obstructive, but individuals appraise themselves capable of coping with it, e.g., being mugged in the street but wanting to defend themselves.
-In the context of CCT, where participants sit in front of a computer, different forms of anger may likely occurred, probably involving an inability to cope.
-The literature suggests that the affective label of interest would be associated with a balanced mix between coping, i.e., the ability to understand, and novelty [41].
-Although our results showed a correlation between interest and coping, an inverse correlation was found between interest and novelty, i.e., higher interest coincided with diminished novelty.
-This could be explained by the fact that the annotations in our study focused on only one sub-dimension of novelty, i.e., predictability, and did not take into account the other existing sub-dimensions such as complexity and unfamiliarity.
+In the context of CCT, where participants sit in front of a computer, different forms of anger may likely occur, probably involving an inability to cope.
 As expected, the affective label of surprise is highly correlated with novelty.
-Even though the surprise affective label is generally considered to be positive or negative depending on the context, it seems to be rather negative in our results as it is correlated with unpleasantness, goal obstructiveness, and no coping.
+Even though the surprise affective label is generally considered to be positive or negative depending on the context, it seems to be rather negative in our results as it is correlated with goal obstructiveness, and no coping.
 This could be explained by the fact that the occurrence of unexpected situations were more related to negative events such as unexpected negative feedback from the virtual assistant.
-Although all the affective labels reported in the Figure 5 seem to be correlated with the appraisal dimensions, with the exception of sad, some correlations actually reflect the result of only one annotator.
-Affective labels were thus defined as irrelevant in the context AI-assisted CCT if no significant correlation is found with any appraisal dimension for at least two annotators, which was the case for several labels, as reported in Table II; labels consistent with the appraisal dimensions are marked with ✓.
 Figure 6: Overview of the predictive experiments performed on the corpus.
@@
 A core set of the likely most relevant affective labels in the context of AI-assisted CCT was defined based on our analyses.
-Four main criteria of selection were used: 1) the frequency is significantly higher than the average on all affective labels, 2) the agreement on presence is significantly higher than chance, 3) the agreement on intensity is sufficiently high, and 4) there is no significant correlation with at least one dimension for at least two annotators.
+Four main criteria of selection were used: (i) the frequency is significantly higher than the average on all affective labels; (ii) the agreement on presence is significantly higher than chance; (iii) the agreement on intensity is sufficiently high; and (iv) there is a significant correlation, in the same direction, with at least one dimension for at least two annotators.
 Ten affective labels meet these four criteria, five positive and five negative, cf.
@@
 In brief, we extract different types of representations (expert or deep) from audio, textual, and visual data, and train specific models for the prediction of the affective labels and dimensions.
-All the experiments are based on a single partitioning of the data set that was divided into three partitions: training, validation and test set, consisting of 56 %, 23 % and 21 % of the participants respectively, c.f.
-Table IV.
-These partitions were defined to ensure that the statistics – 1) ratio of female vs.
-male gender, 2) ratio of senior vs MCI participants, 3) ratio of participants based on affective induction, and 4) distribution of education levels – were preserved in the different population groups.
-All results are reported on the test set.
-In the following, we introduce the used representations of the audio, visual, and textual data, along with the predictive models and strategies, and then report on the results.
-TABLE IV: Summary statistics of the partitions.
+All experiments were conducted using a single partitioning of the dataset, divided into three subsets: training, validation and test, comprising 56 %, 23 % and 21 % of the participants respectively; see Table IV.
+The partitions were carefully designed to minimise bias while maintaining key statistical properties, including: (i) the gender ratio (female vs.
+male); (ii) the proportion of healthy vs.
+MCI participants; (iii) the distribution of participants across affective induction conditions; (iv) the spectrum of education levels; and (v) the distribution of label presence, label intensity, and dimensional intensity.
+Heatmaps depicting these distributions are available in section IV of the supplementary material.
+TABLE IV: Summary statistics of the partitions
 Training Val Test Total
@@
 VI-A Representations of audio, textual and video data
-We included both expert and deep representations of audio, visual, and textual [^6] In order to make the comparison fair, textual data did not rely on the human transcriptions but on the outputs of an ASR, which is evaluated in details in section IV.A of the supplementary material.
+We included both expert and deep representations of audio, visual, and textual [^4] In order to make the comparison fair, textual data did not rely on the human transcriptions but on the outputs of an ASR, which is evaluated in detail in section IV.A of the supplementary material.
 data to evaluate their respective contribution in the analysis of affect.
-Expert based features rely on the statistical description of perceptually salient patterns in the data, and involve relatively light-weight computational resources, whereas deep representations exploit self-supervised learning (SSL) methods on large scale datasets [45, 46, 47].
+Expert based features rely on the statistical description of perceptually salient patterns in the data, and involve relatively light-weight computational resources, whereas deep representations exploit self-supervised learning (SSL) methods on large scale datasets [35].
 VI-A1 Expert based representations
-We used the Mel-scale Filter Banks (MFBs), Term Frequency - Inverse Document Frequency (TF-IDF), and Facial Action Units (FAUs), as representations of the audio, textual, and visual data, respectively.
+We used the Mel-scale Filter Banks (MFBs) [36], Term Frequency - Inverse Document Frequency (TF-IDF) [37], and Facial Action Units (FAUs) [38], as representations of the audio, textual, and visual data, respectively.
 MFBs are popular features in acoustic processing.
@@
 Deep representations of audio, textual, and video data include the following models: Wav2Vec2 [41], BERT [42], and CLIP [43], respectively.
-We used a multilingual Wav2Vec2 model that exploits both convolutional and transformer layers to distinguish positive and negative samples of speech in a contrastive approach [46].
+We used a multilingual Wav2Vec2 model that exploits both convolutional and transformer layers to distinguish positive and negative samples of speech in a contrastive approach.
 The model [^5] https://huggingface.co/voidful/wav2vec2-xlsr-multilingual-56 we used has a large architecture, and was trained on 56 different languages [44], including French, and then fine-tuned for ASR.
@@
 Our first experiments revealed that the Linear model consistently outperformed both the MLP and GRU models.
-In order to predict the time-continuous annotations of the dimensions, we had to re-sample some features to match the annotation frame rate, which was 10 Hz.
-Audio representations, which had a sampling rate of 100 Hz for MFBs, and 50 Hz for Wav2vec2, were simply averaged into chunks of 10 ms to match the sampling frequency of the annotations; the same strategy was applied to the FAUs and CLIP visual representations.
-Regarding textual representations, we only exploited BERT as the TF-IDF returns a single feature vector for the whole sequence.
-Textual features were aligned to the size of the sequence using a spline interpolation, which provided a smooth and continuous representation of the features over time.
-As the Linear and MLP models did not achieve good performance for the prediction of time-continuous dimensions, we further experimented with various GRU models:
+To predict the time-continuous annotations of the dimensions, it was necessary to resample certain features to align with the annotation frame rate of 10 Hz.
+Audio representations, sampled at 100 Hz for MFBs, and 50 Hz for Wav2vec2, were averaged into 10 ms chunks to match the annotation frequency.
+The same strategy was applied to the FAUs and CLIP visual representations.
+For textual representations, we utilised BERT exclusively, as TF-IDF provides a single feature vector for the entire sequence.
+Textual features were aligned to the sequence length using spline interpolation, ensuring a smooth and continuous representation of the features over time.
+Since the Linear and MLP models did not perform well in predicting time-continuous dimensions, we explored various GRU model architectures:
   1. 1.
-    One-Layer GRU: A single-layered GRU with the hidden size equal to half of the input feature size, or equal to the input feature size.
+    One-Layer GRU: A single-layered GRU with a hidden size set to either half the input feature size or equal to the input feature size.
   2. 2.
@@
   3. 3.
-    Three-Layers GRU: Similar to before, but with three hidden layers. A set of fixed hidden sizes was also tested: 128, 256, and 512.
-Results showed that a GRU model with three hidden layers of the same size as the input was optimal for deep representations, whereas the same model with a hidden size of 256 yielded the best results for the hand-crafted representations.
-For multimodal fusion, we experimented with middle and late fusion techniques to deal with the asynchronous inputs.
-The best performing approach for both labels and dimension summaries was a decision-based fusion using an Ordinary Least Squares (OLS) regression model [^10] https://scikit-learn.org.
-Fusion of the time-continuous dimensions was operated with a GRU as a weighted averaging layer to take into account the temporal dynamics of the features.
-The weights of the GRU were optimised on the training set.
+    Three-Layers GRU: Similar to before, but with three hidden layers.
+Additionally, we tested a fixed set of hidden sizes across all GRU models: 128, 256, and 512.
+The results indicated that a GRU model with three hidden layers, each matching the input size, performed optimally for deep representations.
+In contrast, the model with a hidden size of 256 produced the best results for hand-crafted representations.
+For multimodal fusion, we explored both middle and late fusion techniques to address the issue of asynchronous inputs.
+The best performing approach for both labels and dimension summaries was decision-based fusion using an Ordinary Least Squares (OLS) regression model [^8] https://scikit-learn.org.
+For time-continuous dimensions, fusion was achieved through a GRU functioning as a weighted averaging layer, effectively capturing the temporal dynamics of the features.
+The GRU weights were optimised on the training set.
 All the models were trained with the Adam optimizer and an initial learning rate of $10^{-3}$ for affective labels and $10^{-4}$ for dimensions.
 The batch size and gradient accumulation were set to 10 for labels and 1 for dimensions.
-The maximum number of training epochs was set to 50 and an early stopping strategy was used with five epochs.Training of the models was done using the Mean Squared Error (MSE) as loss function, and for evaluation, the Concordance Correlation Coefficient (CCC) was calculated on all the sequences [56].
+The maximum number of training epochs was set to 50 and an early stopping strategy was used with five epochs.
+The models were trained using the Mean Squared Error (MSE) as the loss function, and for evaluation, the Concordance Correlation Coefficient (CCC) was calculated on all sequences [47].
 All experiments were carried out using Pytorch [^9] https://pytorch.org[48], with random seeds set to zero.
@@
 VI-C Affect prediction for the core set labels
-Results obtained on the intensity of the labels are reported in Table V and show some strong dependencies between the modalities and the affective labels (see section III.B of the supplementary material for further details on prediction bias based on population demographics).
-For example, the label “happy” was more accurately evaluated by the video modality, the label “interested” by the textual modality, and the labels “frustrated” and “surprised” by the audio modality with deep representations.
-The choice of using either expert or deep features may further lead to important differences in the results according to the modality.
-For example, “frustrated” was detected with a much higher CCC (.327) with the TF-IDF compared to MFB (.151), but the opposite is observed when using the SSL representations.
-On average, deep representations of the textual, audio, and video modalities performed significantly better compared to their hand-crafted representations.
-The fusion of the different modalities provided the best results on average, and for most affective labels.
-An analysis of the fusion weights showed that the relevance of each modality strongly depends on its representation.
-For example, when using expert features, the TF-IDF presents a much greater influence on the results compared to the MFBs and FAUs, with an averaged fusion coefficient of 0.8, compared to 0.1 for audio and video, whereas the contribution of those two modalities outweigh that of text when using deep representations, with averaged coefficients of .28, .35, and .35 for text, audio, and video respectively.
-TABLE V: Core-set label intensity prediction results (CCC).
+The results for label intensity predictions, presented in Table V, reveal some strong dependencies between modalities and affective labels; see section IV.B of the supplementary material for additional details on prediction bias related to population demographics.
+For instance, the label “happy” was most accurately assessed via the video modality, “interested” through the textual modality, and “frustrated” and “surprised” using the audio modality with deep representations.
+The choice between expert-crafted and deep features also introduced notable differences in the results according to the modality.
+For example, the label “frustrated” achieved a significantly higher CCC (.327) with TF-IDF features compared to MFB (.151), whereas the reverse trend was observed when using SSL representations.
+Overall, deep representations across textual, audio, and video modalities consistently outperformed their hand-crafted counterparts on average.
+The fusion of the different modalities yielded the best results on average and for most affective labels.
+An analysis of the fusion weights revealed that the importance of each modality varied significantly based on its representation.
+For instance, when using expert features, TF-IDF had a substantially greater influence on the results, with an averaged fusion coefficient of 0.8, compared to 0.1 for audio (MFBs) and video (FAUs).
+However, with deep representations, the contributions of audio and video surpassed that of text, with average coefficients of .28, .35, and .35 for text, audio, and video respectively.
+TABLE V: Core-set label intensity prediction results (CCC)
 Label Expert Representations Deep Representations
@@
 Average .199 .170 .202 .333 .229 .280 .255 .380
-An affective label was defined as present for a given sequence if at least one annotator had reported an intensity value for this label.
-We investigated two modeling options for their prediction: learning a binary classification model for each affective label, or learning a threshold from the regression models.
-As the binary classification task achieved significantly worse results compared to a decision based on the regression models, we therefore chose the latter for the experiments.
-Results are shown in table VI.
-Similar to the previous experiments, there exists some dependencies between the best performing modality and the affective labels, even though the difference is less contrasted.
-Overall, we do not observe a significant difference between the expert and deep representations for the different modalities, excepted for the audio modality, and the fusion improved the performance almost systematically.
+An affective label was considered present in a given sequence if at least one annotator reported an intensity value for that label.
+To predict these labels, we explored two modeling approaches: training a binary classification model for each affective label or deriving a threshold from the regression models.
+Since the binary classification approach yielded significantly poorer results compared to the thresholded-based decision from regression models, we opted for the latter in our experiments.
+The results, presented in Table VI, indicate some dependencies between the best-performing modality and the affective labels, although these differences are less pronounced than in previous experiments.
+Overall, no significant performance gap was observed between expert and deep representations across modalities, with the exception of the audio modality.
+Additionally, fusion consistently enhanced performance across most scenarios.
 VI-D Dimensions summaries prediction
-TABLE VI: Core-set label classification results (%UAR).
+TABLE VI: Core-set label classification results (%UAR)
 Label Expert Representations Deep Representations
@@
 Average 67.2 62.1 61.1 69.6 67.7 67.3 62.7 71.6
-Results are reported in table VII for two different gold standards: the average of annotations and the median (see section III.B of the supplementary material for further details on prediction bias based on population demographics).
-Comparing those two gold standards, we notice that the average of the annotations yielded the best performance across the five dimensions.
-Regarding the dimensions, arousal was best predicted by the audio modality and the two others modalities did not improve the performance during fusion.
-Arousal also achieved the lowest scores compared to the others dimensions, which might be explained by the reasons given in V-C.
-Coping was the best predicted dimension where all modalities performed similarly well and showed a complementarity during fusion.
-Goal conduciveness was best predicted by the textual representations, and fusion with the audio and video modalities also improved performance.
-Intrinsic pleasantness was best captured by the video modality; textual representations were relevant only when computed at the whole level of the sequence with the TF-IDF, and deep representations of audio performed similarly well.
-Novelty appeared to be really challenging for the video modality with extremely low scores, whereas textual representations captured better this dimension.
-By analysing the relative importance of each modality based on the coefficients obtained from the linear regression model, we observed that a very low importance was given to audio features for arousal, in particular when using expert features.
-This may explain the decrease in performance for the multimodal approach, compared to using the audio modality alone.
-For the others dimensions, it was observed that the contributions of the modalities were influenced by the complexity of the representations.
-When using deep features, the audio modality emerged as the most influential, but this contribution was reduced for expert features, with textual information showing more important contributions.
-TABLE VII: Dimension summary prediction (CCC).
+Results are presented in table VII for two different gold standards: the average of annotations and the median; see section IV.B of the supplementary material for further details on prediction bias related to population demographics.
+Comparing these gold standards, the average of annotations consistently yielded the best performance across all five dimensions.
+In terms of individual dimensions, arousal was best predicted by the audio modality, with no significant improvement observed from fusion with the other modalities.
+It also had the lowest scores among the dimensions, which might be explained by the factors discussed in Section V-C.
+Coping was the most accurately predicted dimension, with all modalities performing similarly well and demonstrating complementary effects during fusion.
+Goal conduciveness was best predicted by textual representations, with fusion incorporating audio and video modalities further improving performance.
+Intrinsic pleasantness was most effectively captured by the video modality, with textual representations proving relevant only when computed over the entire sequence using TF-IDF.
+Deep audio representations also performed comparably well.
+Finaly, novelty was particularly challenging for the video modality, which yielded extremely low scores, while textual representations proved to be more effective in capturing this dimension.
+By analysing the relative importance of each modality based on the coefficients from the linear regression model, we found that audio features contributed very little to predicting arousal, especially when using expert features.
+This could explain the drop in performance observed with the multimodal approach, compared to using the audio modality alone.
+For the others dimensions, the contributions of each modality were influenced by the complexity of the representations.
+When using deep features, the audio modality emerged as the most influential.
+However, this contribution was diminished when expert features were used, with textual information playing a more significant role.
+TABLE VII: Dimension summary prediction (CCC)
 Modalities
@@
 VI-E Continuous dimensions prediction
-Results are reported in table VIII for three gold standards: the average and the median of the time-continous annotations, and a version based on the Evaluator Weighted Estimator (EWE), which corresponds to a weighted mean of the annotations using the average of the cross-CCC between each rater’s as weights [58] (see section III.B of the supplementary material for further details on prediction bias based on population demographics).
-Overall, the EWE approach systematically outperformed the others two gold standards on the five dimensions.
-Arousal was the best predicted dimension, but this is due to the low variance observed within this dimension, as explained in the previous section.
-For coping, goal conduciveness and novelty, the Wav2vec2 audio representations performed better than the textual features, differently from what was observed for the summarised dimensions, which might be due to the fact that textual features provide a static representation that may not capture the continuous variations within the data.
-Regarding intrinsic pleasantness, the three modalities performed similarly and proved complementary during fusion, which is consistent with the observations made in the previous section.
-In addition, the multimodal approach consistently improved the performance.
+Results are presented in Table VIII for three gold standards: the average and the median of the time-continous annotations, and a version based on the Evaluator Weighted Estimator (EWE), which calculates a weighted mean of the annotations using the average cross-CCC between each rater’s scores as weights [49]; for further details on prediction bias based on population demographics, see section IV.B of the supplementary material.
+Overall, the EWE approach consistently outperformed the others two gold standards across the five dimensions.
+Arousal was the best predicted dimension, but this was primarily due to the low variance observed within this dimension, as explained in the previous section.
+For coping, goal conduciveness and novelty, Wav2vec2 audio representations outperformed textual features, which contrasts with the findings on the summarised dimensions.
+This discrepancy may be attributed to the fact that textual features provide a static representation, which may not capture the continuous variations within the data.
+Regarding intrinsic pleasantness, the three modalities performed similarly and proved complementary during fusion, which is consistent with the previous section’s findings.
+Additionally, the multimodal approach consistently improved performance.
 An analysis of the fusion weights revealed that when using the mean of annotations as the gold standard, the model evenly distributed importance across all modalities.
-However, when using the median or the EWE gold standards, the audio modality appeared as the most significant, whereas text and video contributed similarly.
-An examination of the weights learned by the GRU during fusion showed that the model assigned equal importance to all modalities when using the mean of annotations as the gold standard.
-However, when considering the median and EWE gold standards, the audio modality emerged as the most influential, while the text and video modalities showed comparable contributions to the overall predictions.
-TABLE VIII: Time-continuous dimension prediction (CCC).
+However, when using the median or EWE gold standards, the audio modality emerged as the most significant, while text and video contributed similarly.
+An examination of the weights learned by the GRU during fusion showed that when using the mean of annotations as the gold standard, the model assigned equal importance to all modalities.
+In contrast, with the median and EWE gold standards, the audio modality became the most influential, while text and video modalities contributed similarly to the overall predictions.
+TABLE VIII: Time-continuous dimension prediction (CCC)
 Modalities
@@
 VII Conclusion
-This paper introduced the THERADIA WoZ corpus, an annotated multimodal corpus on affect in the context of healthcare.
-The corpus relies on current affective science findings from recent forms of appraisal theories, and is made available to the affective computing research community.
-The data consist of 39.5h of multimodal sequences from 52 healthy older participants and 9 MCI participants performing WoZ-assisted CCT.
-Affect inductions were performed on 32 of the healthy older participants to maximise the chance of capturing affective episodes in the data.
-All sequences were fully transcribed and partially annotated based on four appraisal dimensions (i.e., novelty, intrinsic pleasantness, goal conduciveness, and coping) plus the arousal dimension, as well as 23 achievement affect labels.
-Dimensions were annotated both in continuous time and in summarised form, while labels were only annotated in summarised form in terms of presence and intensity.
-Combining the results of different analyses yields a core set of ten affective labels (five positive, five negative), evidenced as highly relevant in the context of AI-assisted CCT.
-Extensive baseline experiments and analyses were then conducted to model affect for both labels and dimensions using three modalities of text, audio and video.
-We also explored the fusion of different modalities for both computationally inexpensive feature extraction techniques and more contextual deep representations.
-Results showed some dependencies between the modalities and the different representations of affect.
-This work provides new insights into the field of affect recognition in healthcare and helps to provide additional data for benchmark comparisons.
-In conclusion, we believe that this multidisciplinary work, combining recent advances in affective computing and psychology, fits perfectly with current research challenges.
-Indeed, we believe that such collaborations will benefit both affective computing and psychology by pushing back the boundaries of our knowledge of affect and how it can be modelled.
+This paper presents the THERADIA WoZ corpus, a valuable multimodal dataset for affective computing research in healthcare.
+Based on current affective science and appraisal theories, the corpus includes 39.5 hours of multimodal data from 52 healthy older participants and 9 with MCI performing WoZ-assisted cognitive training.
+To capture affective episodes, affect inductions were applied to 32 healthy older participants.
+The data is fully transcribed and partially annotated across four appraisal dimensions – novelty, intrinsic pleasantness, goal conduciveness, and coping – as well as arousal, and with 23 achievement-related affect labels.
+These dimensions, provided in both in continuous time and summarised forms, highlights ten core affective labels of significant relevance for AI-assisted CCT applications.
+In-depth baseline experiments were conducted to model affect using text, audio, and video modalities, including investigations of multimodal fusion through both computationally efficient feature and contextual deep learning representations.
+Results revealed key dependencies between modalities and affect representations.
+This work provides new insights into the field of affect recognition, offering novel insights into affect recognition in healthcare.
+This work contributes to the ongoing development of affective computing in interaction – and specifically for healthcare – providing a valuable dataset for benchmark comparisons.
+We believe that combining the strengths of affective computing and psychology, as demonstrated in this multidisciplinary study, will significantly advance our understanding of affect and its modeling, benefiting both fields and their applications.
 VIII Acknowledgments